### v1

# On the Maximization of Real Sequences

Abstract
In this paper, we study a maximization problem on real sequences. More precisely, for a given sequence, we are interesting in computing the supremum of the sequence and an index for which the associated term is maximal. We propose a general methodology to solve this maximization problem. The method is based on upper approximations constructed from pairs of eventually decreasing sequences of strictly increasing continuous functions on $[0,1]$ and of scalars in $(0,1)$. Then, we can associate integers with those pairs using inverses on $[0,1]$ of the functions. We prove that such pairs always exist and one provides the index maximizer. In general, such pairs provide an upper bound of the greatest maximizer of the sequence. Finally, we apply the methodology on concrete examples including famous sequences such as logistic, Fibonacci and Syracuse sequences. We also apply our techniques to norm based peaks computation problems on discrete-time linear systems.
1 Introduction
In this paper, denoting by $\mathbb{R}^{\mathbb{N}}$, the set of real sequences, we consider the following computational problem.
Problem 1.
For an element $u\in\mathbb{R}^{\mathbb{N}}$, compute:
  • if it is finite
    $\sup_{n\in\mathbb{N}}u_{n}$
  • and if it exists
    $k\in\operatorname{Argmax}(u):=\{n\in\mathbb{N}:u_{k}=\sup_{m\in\mathbb{N}}u_{m}\}$
The objective of the paper is to propose a theoretical approach explaining how we can compute a supremum and an integer maximizer of a real sequence. The main idea developed in the paper is the construction a formula providing a finite upper bound of the greatest maximizer of the analyzed sequence. This upper bound allows to solve Problem 1 in finite time. Indeed, to compute the supremum of the sequence and to obtain a maximizer, it suffices to compare all the terms from the initial index until the integer part of the upper bound. To limit the number of comparisons, we have to compute the smallest possible upper bound of the greatest maximizer of the sequence.
The study of Problem 1 is motivated by the study of generalized peaks computation problems arising in the verification of dynamical systems (see e.g. [AGM15] in a discrete-time context and for the continuous-time case [MHS20]). Problem 1 can be viewed as a generalization of  [Adj21, Adj25]. A generalized peaks computation problem requires two inputs: a dynamical system and an objective function. Generalized peaks computation problems encompass several verification problems for dynamical systems. The objective function modelizes the quantity of interest for which the evolution has to be analyzed. For example, it can come from energy function definitions [WM07], functional representations of a safe set [AGM15, MHS20] or norms [Shc17, APS18]. Speaking about norms, we can find in some references, a problem dealing with the computation of an upper bound on the sequences of the norms of the powers of a given matrix (see e.g. [Dow13]). The latter work is not directly motivated by dynamical system verification puroposes. Most of the works about norms based generalized peaks computation problems deal with linear systems [PSS18, Shc24]. Moreover, the authors only propose lower and upper bounds over the optimal value of the associated optimization problem. Finally, we find a similar formulation in [AG25]. In [AG25], the generalized peaks computation problem is called a robust-to-dynamics optimization problem. Moreover, the formulation of the authors has an additional constraint called an invariance constraint that forces the vector of the state variable to stay in some invariant set.
In this paper, we aboard maximization problems for sequences and thus the method proposed can be directly applied to generalized peaks computation problems in a discrete-time systems context. The dynamical system input of a generalized peaks computation problem is an autonomous time-invariant discrete-time system characterized by a set of initial conditions $X^{\mathrm{in}}$ and a function $T$ that updates the vector of state variables. If we suppose that the size of the vector of state variables is $d$, then we have $X^{\mathrm{in}}\subset\mathbb{R}^{d}$ and $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$. On the other hand, the objective function is a function of the state space i.e. a function $\varphi:\mathbb{R}^{d}\to\mathbb{R}$. A generalized peaks computation problem consists in maximizing $\varphi$ over the reachable values set generated from $X^{\mathrm{in}}$ and $T$. More formally, a generalized peaks computation problem in a discrete-time context is defined from $X^{\mathrm{in}}$, $T$ and $\varphi$ and aims at computing:
$\sup_{k\in\mathbb{N}}\sup_{x\in X^{\mathrm{in}}}\varphi(T^{k}(x))$ (1)
and, if there exists, an integer $n$ such that
$\sup_{x\in X^{\mathrm{in}}}\varphi(T^{n}(x))=\sup_{k\in\mathbb{N}}\sup_{x\in X^{\mathrm{in}}}\varphi(T^{k}(x)).$
Then writing
$\forall\,k\in\mathbb{N},\ \nu_{k}:=\sup_{x\in X^{\mathrm{in}}}\varphi(T^{k}(x))\text{ and }\nu_{\rm opt}:=\sup_{k\in\mathbb{N}}\nu_{k},$
we obtain a problem of the form of Problem 1.
Problem 1 presented at this level of generality does not appear in the literature. However, the problem of the search of the maximal term of a sequence also appears as an auxiliary problem for studying some famous sequences. The terminology maximum excursion value is used for the famous Syracuse sequences (e.g. [S+99] or [Lag23, PartIII]) or in [PP25] for Juggler sequences [Pic92]. In both cases, this value represents the maximal value of a sequence. Unfortunately, the index achieving this maximal value (if it exists) is not considered. The notion of stopping times well-developed in the context in Syracuse sequences and its generalizations [Lag23, PartIII]) is used in the paper to safely truncate the infinite sequence of maximization problems exposed in Equation (1) to a finite sequence of them.
The main contribution is to propose a general abstract method to solve Problem 1. We develop theoretical algorithms to solve Problem 1 in finite time. To do so, we construct an upper bound of the greatest integer of $\operatorname{Argmax}(u)$. This upper bound is computed from a formula relying on upper bounds of the analyze sequence. Those upper bounds are sequences constructed from a sequence $h:=(h_{k})_{k\in\mathbb{N}}$ of strictly increasing and continuous functions on $[0,1]$ and a sequence $\beta=(\beta_{k})_{k\in;\mathbb{N}}$ of elements in $(0,1)$. The pair $(h,\beta)$ must satisfy, for all $k\in\mathbb{N}$:
$u_{k}\leq h_{k}(\beta_{k}^{k}).$
As for all $k\in\mathbb{N}$, $h_{k}$ is strictly increasing and continuous on $[0,1]$, they are invertible once restricted to $[0,1]$ and we can define, for all $k\in$ such that $h_{k}(0)<u_{k}$:
$\mathfrak{F}_{u}(k,h,\beta)=\dfrac{\ln(h^{-1}(u_{k}))}{\ln(\beta_{k})}.$
In the paper, we prove that the interesting class of upper bounds consists in eventually decreasing elements $(h,\beta)$ for which the set $\{k\in\mathbb{N}:h_{k}(0)<u_{k}\}\cap[N,+\infty)$ is nonempty, $N$ being the smallest integer from which $(h,\beta)$ is decreasing. We prove that such a pair exists if and only if there exists a term of $u$ greater than the limit superior of $u$. Moreover, to use eventually decreasing elements $(h,\beta)$ also guarantees that $\mathfrak{F}_{u}(k,h,\beta)$ is an upper bound of the greatest integer of $\operatorname{Argmax}(u)$ for all $k\geq N$. We then apply our techniques to well-known sequences: the ratio of two consecutive terms of a Fibonacci sequence; logistic sequences; Syracuse sequences and the norm based generalized peaks computation problem in the case of discrete-time linear dynamics. We apply the technique developed in [Adj25] using quadratic Lyapunov functions and discuss how to use the sequential approach developed in this paper. Contrarily to [Dow13, PS16, PSS18, Shc24], where lower and upper bounds on the value $\nu_{\rm opt}$ are only proposed, we compute the exact value of $\nu_{\rm opt}$.
The paper is organized as follows. Section 2 is devoted to preliminary results on the supremum of real sequences. In Section 3, we develop the main results of the paper. We give the framework and useful sets of upper bounds over the analyzed sequence. We also prove that those sets are nonempty in the interesting cases. We present the main properties of the function $\ Fu$. We end Section 3 by presenting theoretical algorithms to solve Problem 1. Section 4 describes some applications of our techniques to concrete sequences. This includes an application to Fibonacci sequences, logistic sequences, Syracuse sequences and to the peak computation problem presented in [APS18].
2 Preliminaries on the supremum of a real sequence
We provide useful notations concerning the supremum and the limit superior of a sequence.
Notation 1.
Let $u\in\mathbb{R}^{\mathbb{N}}$. We write $\mathfrak{s}(u):=\displaystyle{\sup_{n\in\mathbb{N}}u_{n}}$ and $\boldsymbol{\ell}(u):=\displaystyle{\limsup_{n\to+\infty}u_{n}}$.
In this paper, we are interested in sequences for which the supremum is finite and the set containing such sequences is denoted by $\Lambda$.
Notation 2.
$\Lambda:=\{u\in\mathbb{R}^{\mathbb{N}}:\mathfrak{s}(u)<+\infty\}$.
Proposition 1.
We have the asymptotic characterization of elements of $\Lambda$:
$u\in\Lambda\iff\boldsymbol{\ell}(u)\in\mathbb{R}\cup\{-\infty\}.$
Proof.
The direct implication follows from $\mathfrak{s}(u)\geq\boldsymbol{\ell}(u)$. Now assume that $\boldsymbol{\ell}(u)\in\mathbb{R}\cup\{-\infty\}$. If $\boldsymbol{\ell}(u)=-\infty$, then for some $N\in\mathbb{N}$, we have $u_{n}\leq u_{0}$ for all $n\geq N$ and $\mathfrak{s}(u)=\max_{0\leq k\leq N-1}u_{k}\in\mathbb{R}$. Now suppose that $\boldsymbol{\ell}(u)\in\mathbb{R}$ and let $n\in\mathbb{N}$ such that $\sup_{k\geq n}u_{k}\leq\boldsymbol{\ell}(u)+1$. Then, $\mathfrak{s}(u)\leq\displaystyle{\max\{\max_{0\leq j\leq n-1}u_{j},\boldsymbol{\ell}(u)+1\}}\in\mathbb{R}$. ∎
This asymptotic characterization of elements of $\Lambda$ leads to consider new sets of integers dealing with the tails of a sequence $u$:
$\Delta_{u}:=\left\{k\in\mathbb{N}:\max_{0\leq j\leq k}u_{j}\geq\sup_{j>k}u_{j}\right\}\ \text{ and }\ \Delta^{s}_{u}:=\left\{k\in\mathbb{N}:\max_{0\leq j\leq k}u_{j}>\sup_{j>k}u_{j}\right\}$
and their respective infimum:
$\mathbf{K}_{u}=\inf\Delta_{u}\quad\text{ and }\quad\mathbf{K}^{s}_{u}=\inf\Delta^{s}_{u}.$
We recall that $\mathbf{K}_{u}<+\infty$ (resp. $\mathbf{K}^{s}_{u}<+\infty$) if and only if $\Delta_{u}\neq\emptyset$ (resp. $\Delta^{s}_{u}\neq\emptyset$). Furthermore, if the sets of integers defined above are nonempty, they contain their infimum.
Finally, we will also need to check whether a term of the analyzed sequence is greater than its limit superior and thus we introduce the following sets of integers:
${\mathrm{R}}_{u}:=\left\{k\in\mathbb{N}:u_{k}\geq\boldsymbol{\ell}(u)\right\}\quad\text{ and }\quad{\mathrm{R}}_{u}^{s}:=\left\{k\in\mathbb{N}:u_{k}>\boldsymbol{\ell}(u)\right\}.$
Proposition 2.
If $u\notin\Lambda$ then $\Delta_{u}=\emptyset$ and ${\mathrm{R}}_{u}=\emptyset$. Obviously, it also implies that $\Delta^{s}_{u}=\emptyset$ and ${\mathrm{R}}_{u}^{s}=\emptyset$.
Proof.
If $u\notin\Lambda$ then $\boldsymbol{\ell}(u)=+\infty$ and as $u\in\mathbb{R}^{\mathbb{N}}$, we get ${\mathrm{R}}_{u}=\emptyset$. As $\boldsymbol{\ell}(u)=+\infty$, for all $n\in\mathbb{N}$, we can find $N>n$ such that $\sup_{k>n}u_{k}\geq\sup_{k\geq N}u_{k}\geq 1+\max_{0\leq l\leq n}u_{l}$ and $\Delta_{u}=\emptyset$. ∎
Proposition 3.
Let $u\in\Lambda$. The following assertions hold:
  1. 1.
    $\Delta^{s}_{u}\neq\emptyset\iff{\mathrm{R}}_{u}^{s}\neq\emptyset$;
  2. 2.
    $\Delta_{u}\neq\emptyset\iff{\mathrm{R}}_{u}\neq\emptyset$;
  3. 3.
    $\Delta^{s}_{u}=\emptyset\iff\mathfrak{s}(u)=\boldsymbol{\ell}(u)$.
Proof.
1. Suppose that $\Delta^{s}_{u}\neq\emptyset$ and let $k\in\Delta^{s}_{u}$ then we have for some $\ell\leq k$, $u_{\ell}=\max_{0\leq j\leq k}u_{j}>\sup_{j>k}u_{j}\geq\boldsymbol{\ell}(u)$ and $\ell\in{\mathrm{R}}_{u}^{s}\neq\emptyset$. Now assume that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. If $\boldsymbol{\ell}(u)=-\infty$ meaning that ${\mathrm{R}}_{u}^{s}=\mathbb{N}$, for some $N\in\mathbb{N}^{*}$ $\sup_{0\leq k\leq N}u_{k}\geq u_{0}>u_{0}-1\geq\sup_{k>N}u_{k}$ and $N\in\Delta^{s}_{u}$. Now assume that $\boldsymbol{\ell}(u)\in\mathbb{R}$ and suppose that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. Let $k\in{\mathrm{R}}_{u}^{s}$ and $\varepsilon=u_{k}-\boldsymbol{\ell}(u)>0$. Let $n\in\mathbb{N}$, such that $\sup_{j>n}u_{j}\leq\boldsymbol{\ell}(u)+\varepsilon/2=\frac{u_{k}+\boldsymbol{\ell}(u)}{2}<u_{k}$. Then $k\leq n$, $\max_{0\leq j\leq n}u_{j}\geq u_{k}>\sup_{j>n}u_{j}$ and $n\in\Delta^{s}_{u}\neq\emptyset$.
2. The implication $\Delta_{u}\neq\emptyset\implies{\mathrm{R}}_{u}\neq\emptyset$ can be proved as in 1. Now, suppose that ${\mathrm{R}}_{u}\neq\emptyset$. If ${\mathrm{R}}_{u}^{s}\neq\emptyset$, the result follows from 1. So, we assume that ${\mathrm{R}}_{u}^{s}=\emptyset$. Let $k\in{\mathrm{R}}_{u}$. As ${\mathrm{R}}_{u}^{s}=\emptyset$, we have $u_{k}=\boldsymbol{\ell}(u)$ and for all $j\in\mathbb{N}$, $u_{j}\leq\boldsymbol{\ell}(u)=u_{k}$. Then $u_{k}=\mathfrak{s}(u)$ and $\max_{0\leq j\leq k}u_{j}\geq\sup_{j>k}u_{j}$ and $k\in\Delta_{u}\neq\emptyset$.
3. From 1. we have $\Delta^{s}_{u}=\emptyset\iff{\mathrm{R}}_{u}^{s}=\emptyset$. As $\Delta^{s}_{u}=\emptyset$ iff $\forall\,k\in\mathbb{N},\ u_{k}\leq\boldsymbol{\ell}(u)$ iff $\mathfrak{s}(u)\leq\boldsymbol{\ell}(u)$ and since $\mathfrak{s}(u)\geq\boldsymbol{\ell}(u)$ is always true, we conclude that $\mathfrak{s}(u)=\boldsymbol{\ell}(u)$. ∎
Proposition 4.
Let $u\in\Lambda$. The following assertions hold:
  1. 1.
    $\operatorname{Argmax}(u)\subseteq\Delta_{u}$;
  2. 2.
    $\operatorname{Argmax}(u)\neq\emptyset\iff\Delta_{u}\neq\emptyset$;
  3. 3.
    $\mathbf{K}_{u}=\inf\operatorname{Argmax}(u)$;
  4. 4.
    $\operatorname{Argmax}(u)=\emptyset$ or $\operatorname{Argmax}(u)$ unbounded $\implies\mathfrak{s}(u)=\boldsymbol{\ell}(u)$;
  5. 5.
    If $\Delta^{s}_{u}\neq\emptyset$ then $\mathbf{K}^{s}_{u}=\max\operatorname{Argmax}(u)$.
Proof.
1. Let $k\in\operatorname{Argmax}(u)$. We have $u_{k}=\mathfrak{s}(u)$; thus $\max_{0\leq j\leq k}u_{j}=u_{k}\geq\sup_{j>k}u_{j}$ and $k\in\Delta_{u}$.
2. From 1. $\operatorname{Argmax}(u)\neq\emptyset\implies\Delta_{u}\neq\emptyset$. Now as for all $k\in\Delta_{u}$, $\mathfrak{s}(u)=\max_{0\leq j\leq k}u_{j}$, there exists some $j\in\operatorname{Argmax}(u)\cap[0,k]$. We conclude that $\Delta_{u}\neq\emptyset\implies\operatorname{Argmax}(u)\neq\emptyset$.
3. Assume that $\mathbf{K}_{u}<+\infty$. From 1., $\mathbf{K}_{u}\leq k$ for all $k\in\operatorname{Argmax}(u)$. Moreover, as $\mathbf{K}_{u}\in\Delta_{u}$, $\mathfrak{s}(u)=\max_{0\leq j\leq\mathbf{K}_{u}}u_{j}$. Then, there exists $j\in\operatorname{Argmax}(u)\cap[0,\mathbf{K}_{u}]$. From 1., we also have $\mathbf{K}_{u}\leq j$. Finally, $j=\mathbf{K}_{u}$ and $\mathbf{K}_{u}\in\operatorname{Argmax}(u)$. If $\mathbf{K}_{u}=+\infty$, then $\Delta_{u}=\emptyset$ and, from 2., this is equivalent to $\operatorname{Argmax}(u)=\emptyset$ i.e. $\min\operatorname{Argmax}(u)=+\infty$.
4. Assume that $\operatorname{Argmax}(u)=\emptyset$. Then $\Delta_{u}=\emptyset$ and so $\Delta^{s}_{u}=\emptyset$. We conclude from the third statement of Prop. 3. Now if $\operatorname{Argmax}(u)$ is nonempty but unbounded, then for all $k\in\Delta_{u}$, there exists $j>k$ such that $j\in\operatorname{Argmax}(u)$ and thus $\max_{0\leq l\leq k}u_{l}=\max_{l>k}u_{l}$ and so for all $k\in\Delta_{u}$, $k\notin\Delta^{s}_{u}$ and $\Delta^{s}_{u}=\emptyset$. We conclude again using the third statement of Prop. 3.
5. Assume that $\Delta^{s}_{u}\neq\emptyset$. Then $\operatorname{Argmax}(u)$ is nonempty and let $k\in\operatorname{Argmax}(u)$. We have $k\leq\mathbf{K}^{s}_{u}$. Otherwise, as $\mathbf{K}^{s}_{u}\in\Delta^{s}_{u}$, $\mathfrak{s}(u)\geq\max_{0\leq j\leq\mathbf{K}^{s}_{u}}u_{j}>\sup_{j>\mathbf{K}^{s}_{u}}u_{j}=u_{k}=\mathfrak{s}(u)$. Suppose that $\mathbf{K}^{s}_{u}\notin\operatorname{Argmax}(u)$ and let $\overline{l}=\max\{0\leq l\leq\mathbf{K}^{s}_{u}:u_{l}=\sup_{0\leq j\leq\mathbf{K}^{s}_{u}}u_{j}\}$. As $\mathbf{K}^{s}_{u}\in\Delta^{s}_{u}$, $\sup_{l>\mathbf{K}^{s}_{u}}u_{l}<\max_{0\leq l\leq\mathbf{K}^{s}_{u}}u_{l}=u_{\overline{l}}$. By definition of $\overline{l}$, $u_{\overline{l}}>\max_{\overline{l}<l\leq\mathbf{K}^{s}_{u}}u_{l}$. Finally, $u_{\overline{l}}=\max_{0\leq l\leq\overline{l}}u_{l}>\sup_{l>\overline{l}}u_{l}$ and $\overline{l}<\mathbf{K}^{s}_{u}$. This contradicts the minimality of $\mathbf{K}^{s}_{u}$ and $\mathbf{K}^{s}_{u}\in\operatorname{Argmax}(u)$. ∎
Note that $\operatorname{Argmax}(u)$ can be nonempty and bounded when $\Delta^{s}_{u}$ is empty. In this case, we do have $\mathbf{K}^{s}_{u}=+\infty$ and thus $\max\operatorname{Argmax}(u)<\mathbf{K}^{s}_{u}$. Moreover, by the third statement of Prop 3, we have $\mathfrak{s}(u)=\boldsymbol{\ell}(u)$. However, the sequence finitely touches $\boldsymbol{\ell}(u)$ before approaching, up to a subsequence, $\boldsymbol{\ell}(u)$ from below (strictly).
3 Resolution of Problem 1 from an upper approximation of the analyzed sequence
In Section 2, we identify the two integers which characterize the set of maximizer of a real sequence. We propose in this paper, given a real sequence $u$ for which ${\mathrm{R}}_{u}^{s}\neq\emptyset$, to compute an upper bound of the greatest maximizer $\mathbf{K}^{s}_{u}$. Once an upper bound $K$ of $\mathbf{K}^{s}_{u}$ computed, we are able to solve Problem 1 by comparing the terms $u_{0},u_{1},\ldots,u_{K}$. As $K$ is an upper bound of $\mathbf{K}^{s}_{u}$, we know that the optimal value $\mathfrak{s}(u)$ belongs to this finite truncation of the sequence $u$. To avoid useless computations, $K$ must be as close as possible to $\mathbf{K}^{s}_{u}$.
3.1 The set of upper bound sequences and the upper bound functional $\mathfrak{F}_{u}$
Given a real sequence $u$, $\mathbf{K}^{s}_{u}$ is a index of $u$. Besides, to compute an upper bound of $\mathbf{K}^{s}_{u}$ from a closed formula can only be done from the available data i.e. the terms of $u$. Thus, we have to convert terms of $u$ into indices. One class of sequences for which the conversion is easy is the class of convergent geometric sequences. However, to keep the highest possible level of generality, we apply bijections to convergent geometric sequences and those particular sequences are used as upper bounds of $u$. To use such sequences as upper bounds forces us to only consider (strictly) increasing surjection and positive convergent geometric sequences. More formally, upper bound candidates of $u$ are sequences $\omega:=(\omega_{k})_{k\in\mathbb{N}}$ such that for all $k\in\mathbb{N}$, there exists $h_{k}:\mathbb{R}\to\mathbb{R}$ strictly increasing on $[0,1]$ and surjective (as a function from $[0,1]$ to $[h_{k}(0),h_{k}(1)]$) and $\beta_{k}\in(0,1)$ verifying $\omega_{k}=h_{k}(\beta_{k}^{k})$. We can easily convert the terms $\omega_{k}$ into indices $k$. With a suitable choice of sequences of functions $(h_{k})_{k\in\mathbb{N}}$, the obtained indices $k$ will be upper bounds of $\mathbf{K}^{s}_{u}$.
Actually, the upper bound candidate $\omega$ are completely determined by a sequence $h$ of strictly increasing surjections $h_{k}$ on $[0,1]$ and a sequence $\beta$ in $(0,1)$. As a strictly increasing function $g:[a,b]\to[g(a),g(b)]$ is surjective if and only if it is continuous on $[a,b]$, we introduce the set of functions:
$\Omega([0,1]):=\{h:\mathbb{R}\mapsto\mathbb{R}:h\text{ strictly increasing andcontinuous on }[0,1]\}.$
[…]
Then, we use the standard notation $\Omega([0,1])^{\mathbb{N}}$ for the sequences of elements of $\Omega([0,1])$. For $u\in\mathbb{R}^{\mathbb{N}}$, we define the following set of the upper bounds of $u$:
$\Gamma(u):=\left\{(h,\beta)\in\Omega([0,1])^{\mathbb{N}}\times(0,1)^{\mathbb{N}}:u_{k}\leq h_{k}(\beta_{k}^{k}),\ \forall\,k\in\mathbb{N}\right\}.$
We need sometimes to separate functional and scalar parts of an element in $\Gamma(u)$ and we introduce the two following sets:
$\Gamma_{\rm f}(u):=\left\{h\in\Omega([0,1])^{\mathbb{N}}:\exists\,\beta\in(0,1)^{\mathbb{N}}\text{ s.t. }(h,\beta)\in\Gamma(u)\right\}$
and for $h\in\Gamma_{\rm f}(u)$:
$\Gamma_{\rm s}(u,h):=\left\{\beta\in(0,1)^{\mathbb{N}}:(h,\beta)\in\Gamma(u)\right\}.$
In the sequel, we will consider special sequences in $\Gamma(u)$ and $\Gamma_{\rm f}(u)$. First, we define the set $\Gamma_{{\rm ev}}^{\rm cst}(u)$ of eventually constant pairs in $\Gamma(u)$:
$\Gamma_{{\rm ev}}^{\rm cst}(u):=\{(h,\beta)\in\Gamma(u):\exists\,(g,\gamma,N)\in\Omega([0,1])\times(0,1)\times\mathbb{N}\ \forall\,n\geq N,\ (h_{n},\beta_{n})=(g,\gamma)\},$
For $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm cst}(u)$, we denote by $\mathfrak{c}(h,\beta)$ the smallest index from which the sequence $(h,\beta)$ is constant.
We will also consider the set ${\Gamma}_{0}^{\rm cst}(u)$ of constant pairs in $\Gamma(u)$:
${\Gamma}_{0}^{\rm cst}(u):=\{(h,\beta)\in\Gamma(u):\exists\,(g,\gamma)\in\Omega([0,1])\times(0,1)\ \forall\,n\in\mathbb{N},\ (h_{n},\beta_{n})=(g,\gamma)\}.$
For $(h,\beta)\in{\Gamma}_{0}^{\rm cst}(u)$, we have obviously $\mathfrak{c}(h,\beta)=0$.
From elements of $\Gamma(u)$, we can derive a function called the upper bound functional. We will see in that Subsection 3.3, for some selections of elements of $\Gamma(u)$, the function produces upper bounds of $\mathbf{K}^{s}_{u}$.
Definition 1 (Upper bound Functional).
Let $u\in\mathbb{R}^{\mathbb{N}}$. The upper bound functional is the function $\mathfrak{F}_{u}:\mathbb{N}\times\Omega([0,1])^{\mathbb{N}}\times(0,1)^{\mathbb{N}}\mapsto\mathbb{R}_{+}\cup\{+\infty\}$ defined as follows. For all $k\in\mathbb{N}$, $h\in\Omega([0,1])^{\mathbb{N}}$ and $\beta\in(0,1)^{\mathbb{N}}$, we have:
$\mathfrak{F}_{u}(k,h,\beta):=\left\{\begin{array}[]{lr}\dfrac{\ln(h_{k}^{-1}(u_{k}))}{\ln(\beta_{k})}&\text{ if }(h,\beta)\in\Gamma(u)\text{ and }u_{k}>h_{k}(0)\\ +\infty&\text{otherwise}\end{array}\right.$ (2)
To be complete, $\mathfrak{F}_{u}$ is defined and equal to +$\infty$+ ∞ for the elements of $\Omega([0,1])^{\mathbb{N}}\times(0,1)^{\mathbb{N}}$ outside $\Gamma(u)$ and for elements of $\Gamma(u)$ for which $h_{k}^{-1}(u_{k})\leq 0$. Later, we will prove that, for a pertinent subset of $\Gamma(u)$, ${\mathrm{R}}_{u}^{s}$ is nonempty (and thus $\mathbf{K}^{s}_{u}<+\infty$) if and only if we can find an element $(h,\beta)$ of this subset for which there exists $k\in\mathbb{N}$ such that $u_{k}>h_{k}(0)$. Hence, if ${\mathrm{R}}_{u}^{s}$ is empty, $\mathbf{K}^{s}_{u}$ is infinite and $\mathfrak{F}_{u}$ is also infinite on this subset.
The existence of an integer $k$ such that $u_{k}>h_{k}(0)$ plays an important role in the choice of the upper bound candidate. For example, if $u\in\Lambda$, the constant sequence $h$ identically equal to $g:x\to\mathfrak{s}(u)+x$ belongs to ${\Gamma}_{\rm f,0}^{\rm cst}(u)$ but there is no $k\in\mathbb{N}$ such that $u_{k}>h_{k}(0)$ even if ${\mathrm{R}}_{u}^{s}$ is nonempty. This sequence cannot be used to construct an upper bound of $\mathbf{K}^{s}_{u}$ since $\mathfrak{F}_{u}$ evaluated at this $h$ is infinite. This simple situation leads to introduce a notion of usefulness.
Definition 2 (Useful strictly increasing continuous functions).
Let $u\in\mathbb{R}^{\mathbb{N}}$. A function $h\in\Gamma_{\rm f}(u)$ is said to be useful for $u$ if the set :
$\mathcal{S}(u,h):=\{k\in\mathbb{N}:u_{k}>h_{k}(0)\}$ (3)
is nonempty.
By extension, a pair $(h,\beta)\in\Gamma(u)$ is useful for $u$ if $h$ is useful for $u$.
In the sequel, if there is no ambiguity about the sequence, we will simply say useful instead of useful for $u$.
Unfortunately, as we will see in Example 1, usefulness is not enough to compute safely an upper bound of $\mathbf{K}^{s}_{u}$ from $\mathfrak{F}_{u}$.
Example 1.
Let $u\in\mathbb{R}^{\mathbb{N}}$ such that $\mathfrak{s}(u)\in\mathbb{R}_{+}\cup\{+\infty\}$. We define for all $k\in\mathbb{N}$:
$h_{k}:x\mapsto\left(\frac{|u_{k}|+1}{|u_{k}|+0.5}\right)^{k}|u_{k}|x\text{ and}\beta_{k}:=\left(\frac{|u_{k}|+0.5}{|u_{k}|+1}\right).$ (4)
For all $k\in\mathbb{N}$, the functions $h_{k}$ are clearly in $\Omega([0,1])$ and $\beta_{k}\in(0,1)$. Moreover, we have for all $k\in\mathbb{N}$, $u_{k}\leq h_{k}(\beta_{k}^{k})$ and in particular we have $u_{k}=h_{k}(\beta_{k}^{k})$ for all $k\in\mathbb{N}$ such that $u_{k}\geq 0$. Let us define $h:=(h_{k})_{k}$ and $\beta:=(\beta_{k})_{k}$. We have for all $k\in\mathbb{N}$, $h_{k}(0)=0$. As $\mathfrak{s}(u)$ is supposed to be strictly positive, there exists $k\in\mathbb{N}$ such that $u_{k}>0$ and so $\mathcal{S}(u,h)\neq\emptyset$. Hence, for all $k$ such that $u_{k}>0$:
$\mathfrak{F}_{u}(k,h,\beta)=\dfrac{\ln h_{k}^{-1}(u_{k})}{\ln\beta_{k}}=k$
Obviously, we cannot guarantee that $\mathfrak{F}_{u}(k,h,\beta)$ is an upper bound of $\mathbf{K}^{s}_{u}$ for all $k\geq K$ for some $K$ independent to $\mathbf{K}^{s}_{u}$. ∎
In Example 1, we see that we need to restrict the choice of upper bound candidates. Next, we will see that a good choice is to only pick eventually decreasing elements of $\Gamma(u)$.
We use the standard functional order on $\Omega([0,1])$ i.e. for $f,g\in\Omega([0,1])$, $f\leq g\iff f(x)\leq g(x)$ for all $x\in[0,1]$. We also use the standard weak order on $\Omega([0,1])\times(0,1)$ i.e. for all $(f,\beta),(g,\gamma)\in\Omega([0,1])\times(0,1)$, $(f,\beta)\leq(g,\gamma)\iff f\leq g$ and $\beta\leq\gamma$. Then, we introduce for $u\in\mathbb{R}^{\mathbb{N}}$, the set ${\Gamma}_{\rm f,{\rm ev}}^{\rm dec}(u)$ of eventually decreasing of $\Gamma_{\rm f}(u)$ i.e.
${\Gamma}_{\rm f,{\rm ev}}^{\rm dec}(u):=\{h\in\Gamma_{\rm f}(u):\exists\,N\in\mathbb{N}\ \forall\,n\geq N,\ h_{n+1}\leq h_{n}\}.$
For $h\in{\Gamma}_{\rm f,{\rm ev}}^{\rm dec}(u)$, we denote by $\mathfrak{m}(h)$ the smallest index from which $h$ is decreasing.
The set $\Gamma_{{\rm ev}}^{\rm dec}(u)$ contains all eventually decreasing elements of $\Gamma(u)$ i.e.
$\Gamma_{{\rm ev}}^{\rm dec}(u):=\{(h,\beta)\in\Gamma_{\rm f}(u):\exists\,N\in\mathbb{N}\ \forall\,n\geq N,\ (h_{n+1},\beta_{n+1})\leq(h_{n},\beta_{n})\}.$
For $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$, we denote by $\mathfrak{m}(h,\beta)$ the smallest index from which $h$ and $\beta$ are decreasing.
Then, we define ${\Gamma}_{{\rm f},0}^{\rm dec}(u)$ as the set of decreasing elements of $\Gamma_{\rm f}(u)$, formally:
${\Gamma}_{{\rm f},0}^{\rm dec}(u):=\{h\in\Gamma_{\rm f}(u):\forall\,n\in\mathbb{N},\ h_{n+1}\leq h_{n}\},$
and finally, the set $\Gamma_{0}^{\rm dec}(u)$ of decreasing elements of $\Gamma(u)$, formally:
$\Gamma_{0}^{\rm dec}(u):=\{(h,\beta)\in\Gamma_{\rm f}(u):\forall\,n\in\mathbb{N},\ (h_{n+1},\beta_{n+1})\leq(h_{n},\beta_{n})\}.$
It is straightforward to see that ${\Gamma}_{\rm f,{\rm ev}}^{\rm dec}(u)$ and $\Gamma_{{\rm ev}}^{\rm dec}(u)$ are respectively subsets of ${\Gamma}_{{\rm f},0}^{\rm dec}(u)$ and $\Gamma_{0}^{\rm dec}(u)$. Obviously, for all $h\in{\Gamma}_{{\rm f},0}^{\rm dec}(u)$ and all $(g,\gamma)\in\Gamma_{0}^{\rm dec}(u)$, we have $\mathfrak{m}(h)=\mathfrak{m}(g,\gamma)=0$. Moreover, as constant sequences are decreasing, $\Gamma_{{\rm ev}}^{\rm cst}(u)$ and ${\Gamma}_{0}^{\rm cst}(u)$ are respectively subsets of $\Gamma_{{\rm ev}}^{\rm dec}(u)$ and $\Gamma_{0}^{\rm dec}(u)$. Since an eventually constant sequence can be decreasing before being constant, for all $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm cst}(u)$, we have $\mathfrak{m}(h,\beta)\leq\mathfrak{c}(h,\beta)$.
Remark 1.
We insist on the fact that in the definition of being eventually decreasing both $h$ and $\beta$ are decreasing after $\mathfrak{m}(h,\beta)$. Let $n$ be greater than $\mathfrak{m}(h,\beta)$. As the functional order is used for $h$, we have $h_{n+1}(\beta_{n+1}^{n+1})\leq h_{n}(\beta_{n+1}^{n+1})$. Since $h_{n}$ is strictly increasing and $\beta_{n+1}\in(0,1)$, we have $h_{n}(\beta_{n+1}^{n+1})<h_{n}(\beta_{n+1}^{n})$. Finally, as $\beta_{n+1}\leq\beta_{n}$ and $h_{n}$ is increasing, we get $h_{n}(\beta_{n+1}^{n})\leq h_{n}(\beta_{n}^{n})$. We conclude that $h_{n+1}(\beta_{n+1}^{n+1})<h_{n}(\beta_{n})$.
We could ask if it suffices to only require the weaker condition: $h_{n}(\beta_{n+1}^{n})\leq h_{n}(\beta_{n}^{n})$ for all $n\geq\mathfrak{m}(h,\beta)$. However, in the definition of the upper bound functional (Definition 1), $(h_{n})_{n\in\mathbb{N}}$ is decoupled from $(\beta_{n})_{n\in\mathbb{N}}$. The weaker condition would not be sufficient to compare $\mathfrak{F}_{u}(k,h,\beta)$, $\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)$ and $\mathbf{K}^{s}_{u}$ for $k\leq\mathbf{K}^{s}_{u}$.
We extend usefulness to eventually decreasing elements. Now we require that $\mathcal{S}(u,h)$ contains an integer greater than $\mathfrak{m}(h)$.
Definition 3 (Usefully decreasing).
Let $u\in\mathbb{R}^{\mathbb{N}}$. An element $h\in{\Gamma}_{\rm f,{\rm ev}}^{\rm dec}(u)$ (resp. $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$) is usefully decreasing if and only if $\mathcal{S}(u,h)\cap[\mathfrak{m}(h),+\infty)\neq\emptyset$ (resp. $\mathcal{S}(u,h)\cap[\mathfrak{m}(h,\beta),+\infty)\neq\emptyset$).
An element $h\in\Gamma_{0}^{\rm dec}(u)$ (or $(h,\beta)\in\Gamma_{0}^{\rm dec}(u)$) is usefully decreasing if and only if it is useful in the sense of Definition 2.
We can always construct a decreasing pair from an eventually decreasing one. Later, we prove that the procedure may increase the value of $\mathfrak{F}_{u}$.
Proposition 5.
Let $u\in\mathbb{R}^{\mathbb{N}}$. The following equivalence holds :
$\Gamma_{{\rm ev}}^{\rm dec}(u)\neq\emptyset\iff\Gamma_{0}^{\rm dec}(u)\neq\emptyset.$
Moreover, there exists $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$ usefully decreasing if and only if there exists $(g,\gamma)\in\Gamma_{0}^{\rm dec}(u)$ useful.
Proof.
If $(h,\beta)$ is decreasing then it is eventually decreasing. Now suppose there exists $(g,\gamma)\in\Gamma(u)$ eventually decreasing. Let us write $N=\mathfrak{m}(g,\gamma)$. Then we can define for all integer $n\leq N$, $h_{n}=\max_{i=0,\ldots,N}g_{i}$ and $h_{n}=g_{n}$ for all integer $n>N$. Similarly, we define, for all $n\leq N$, $\beta_{n}=\max_{i=0,\ldots,N}\alpha_{i}$ and for all $n>N$, $\beta_{n}=\alpha_{n}$. The sequences $\beta$ and $h$ are clearly decreasing.
Clearly, if there exists $(g,\gamma)\in\Gamma_{0}^{\rm dec}(u)$ useful then it is also eventually decreasing and usefully decreasing. Now suppose that there exists $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$ usefully decreasing, the construction of $(g,\gamma)$ from above is useful as for some $n\geq\mathfrak{m}(h,\beta)$, $u_{n}>h_{n}(0)=g_{n}(0)$. ∎
3.2 Existence of useful upper bound sequences
We have presented the important sets of upper bound sequences and their basic properties. Before developing the main properties of the upper bound functional $\mathfrak{F}_{u}$ defined in Definition 1, we focus on the existence of useful elements in ${\Gamma}_{0}^{\rm cst}(u)$ and usefully decreasing elements in $\Gamma_{{\rm ev}}^{\rm dec}(u)$ when ${\mathrm{R}}_{u}^{s}$ is nonempty.
First, the set $\Gamma(u)$ is still nonempty. Second, it is easy to see that for all $h\in\Gamma_{\rm f}(u)$, for all $k\in\mathbb{N}^{*}$, $u_{k}<h_{k}(1)$ and $u_{0}\leq h_{0}(1)$. From these two inequalities, we can characterize $\Lambda$ from $\Gamma_{\rm f}(u)$ and describe a simple situation where $u_{0}$ is the maximum of the sequence terms.
Proposition 6.
Let $u\in\mathbb{R}^{\mathbb{N}}$. The following statements hold:
  1. 1.
    $\Gamma(u)\neq\emptyset$.
  2. 2.
    $u\in\Lambda$ if and only there exists $h\in\Gamma_{\rm f}(u)$ such that the set $\{h_{k}(1):k\in\mathbb{N}\}$ is bounded from above.
  3. 3.
    if there exists $h\in{\Gamma}_{{\rm f},0}^{\rm dec}(u)$ and $u_{0}=h_{0}(1)$ then $u\in\Lambda$ and $u_{k}<u_{0}$ for all $k\in\mathbb{N}$.
Proof.
1. The pair $(h,\beta)$ proposed in Equation (4) belongs to $\Gamma(u)$ even if $u\notin\Lambda$ and $\mathfrak{s}(u)$ is not strictly positive.
2. Let us take $h\in\Gamma(u)$ such that $\{h_{k}(1):k\in\mathbb{N}\}$ is bounded from above then $u_{k}\leq h_{k}(\beta_{k}^{k})<h_{k}(1)$ which implies that $\mathfrak{s}(u)\leq\sup\{h_{k}(1):k\in\mathbb{N}\}<+\infty$. On the contrary, if $u\in\Lambda$, then the constant sequence $h$ defined from $f:x\mapsto x+\mathfrak{s}(u)$ belongs to $\Gamma_{\rm f}(u)$ and the set $\{h_{k}(1):k\in\mathbb{N}\}=\{\mathfrak{s}(u)+1\}$.
3. Assume that there exists $h\in\Gamma_{\rm f}(u)$ decreasing such that $u_{0}=h_{0}(1)$. Then for some $\beta\in(0,1)^{\mathbb{N}}$, for all $k\in\mathbb{N}^{*}$, $u_{k}\leq h_{k}(\beta_{k}^{k})<h_{k}(1)\leq h_{0}(1)=u_{0}$. ∎
Remark 2.
The third statement of Proposition 6 implies, if it holds, that the term $u_{0}$ is the supremum of the analyzed sequence $u$. Moreover, the index 0 is the only maximizer of the sequence.
We can go further than the results of Proposition 6 to completely characterize $\Lambda$ from eventually decreasing upper bounds.
Proposition 7.
Let $u\in\mathbb{R}^{\mathbb{N}}$. Then $u\in\Lambda\iff{\Gamma}_{\rm f,{\rm ev}}^{\rm dec}(u)\neq\emptyset$.
Proof.
Assume that $u\in\Lambda$. The constant sequence function of the proof of the second statement of Proposition 6 belongs to ${\Gamma}_{\rm f,{\rm ev}}^{\rm dec}(u)$. Now suppose that $u\notin\Lambda$ and ${\Gamma}_{\rm f,{\rm ev}}^{\rm dec}(u)\neq\emptyset$. Using the second statement of Proposition 6, for all $h\in\Gamma_{\rm f}(u)$, the set $\{h_{k}(1):k\in\mathbb{N}\}$ is unbounded from above. Taking $g\in{\Gamma}_{\rm f,{\rm ev}}^{\rm dec}(u)$, we have for all $k\in\mathbb{N}$, $g_{k}(1)\leq\max\{g_{i}(1):i=0,\ldots,\mathfrak{m}(g)\}\in\mathbb{R}$. This contradicts the second statement of Proposition 6 and thus ${\Gamma}_{\rm f,{\rm ev}}^{\rm dec}(u)$ is empty. ∎
Theorem 1 above is the main result of the section. First, it justifies the finiteness of the upper bound functional $\mathfrak{F}_{u}$ for $u\in\Lambda$ with ${\mathrm{R}}_{u}^{s}\neq\emptyset$. In a second time, it serves to construct constant and non constant elements in $\Gamma(u)$. Finally, the existence of a useful and optimal function is important for the optimality of the proposed function $\mathfrak{F}_{u}$.
Theorem 1 (Existence of a useful optimal affine function).
Let $u\in\Lambda$ such that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. Then, there exist $a>0$, $b\in(0,1)$ and $c\in\mathbb{R}$ such that:
  1. 1.
    $\{k\in\mathbb{N}:u_{k}>c\}\neq\emptyset$;
  2. 2.
    $u_{k}\leq ab^{k}+c$ for all $k\in\mathbb{N}$;
  3. 3.
    $u_{\mathbf{K}^{s}_{u}}=ab^{\mathbf{K}^{s}_{u}}+c$.
Proof.
If ${\mathrm{R}}_{u}^{s}$ is nonempty then $u_{\mathbf{K}^{s}_{u}}=\mathfrak{s}(u)>\boldsymbol{\ell}(u)\in\mathbb{R}\cup\{-\infty\}$ and for all $t\in(\boldsymbol{\ell}(u),\mathfrak{s}(u))$ there exists $k\in\mathbb{N}$, such that $u_{k}>t$. So let $c$ be in $(\boldsymbol{\ell}(u),\mathfrak{s}(u))$. Now as $c>\boldsymbol{\ell}(u)$, there exists $N_{c}\in\mathbb{N}$ s.t. $\sup_{k\geq N_{c}}x_{k}\leq c$. This is the same as for all $k\geq N_{c}$, $u_{k}\leq c$. We conclude that for all $k\geq N_{c}$, $u_{k}\leq c+ab^{k}$ for all $a>0$ and all $b\in(0,1)$. Note that as $c<\mathfrak{s}(u)$ and $\mathbf{K}^{s}_{u}$ is the greatest element of the argmax of $u$ then $N_{c}$ is strictly greater than $\mathbf{K}^{s}_{u}$. Now let introduce the set :
${AB}_{c}:=\left\{(a,b)\in\mathbb{R}_{+}^{*}\times(0,1):u_{\mathbf{K}^{s}_{u}}=\mathfrak{s}(u)=ab^{\mathbf{K}^{s}_{u}}+c\right\}.$
It is easy to see that $S_{c}$ is nonempty, indeed, for all $b$ in $(0,1)$, $(\frac{u_{\mathbf{K}^{s}_{u}}-c}{b^{\mathbf{K}^{s}_{u}}},b)\in S_{c}$. Moreover, if $\mathbf{K}^{s}_{u}>0$, for all $(a,b)\in{AB}_{c}$, for all $0<k<\mathbf{K}^{s}_{u}$, $u_{k}\leq u_{\mathbf{K}^{s}_{u}}=ab^{\mathbf{K}^{s}_{u}}+c<ab^{k}+c$ as $b\in(0,1)$. We also have $u_{0}\leq u_{\mathbf{K}^{s}_{u}}=ab^{\mathbf{K}^{s}_{u}}+c\leq a+c$ as $b\in(0,1)$. Hence, the result $u_{k}\leq ab^{k}+c$ holds on the sets $[0,\mathbf{K}^{s}_{u}]\cap\mathbb{N}$ and $[N_{c},+\infty)\cap\mathbb{N}$ with for every $(a,b)\in S_{c}$. It suffices to find $(a,b)\in S_{c}$, such that $u_{k}\leq ab^{k}+c$ for all $k\in(\mathbf{K}^{s}_{u},N_{c}]\cap\mathbb{N}$. Recall that we have $\mathbf{K}^{s}_{u}+1\leq N_{c}$.
Let us write $(-\infty,0)\ni\gamma:=\max_{k\in(\mathbf{K}^{s}_{u},N_{c}]}(u_{\mathbf{K}^{s}_{u}}-u_{k})/(\mathbf{K}^{s}_{u}-k)$. This cannot be zero since $\mathbf{K}^{s}_{u}$ is the maximal integer $k$ such that $u_{k}=\max_{n\in\mathbb{N}}u_{n}$. Now, let $(a,b)\in\mathbb{R}_{+}\times(0,1)$ and define $f_{a,b}:\mathbb{R}\ni y\mapsto ab^{y}+c$. The function $f_{a,b}$ is derivable and strictly convex and $f^{\prime}_{a,b}(y)=ab^{y}\ln(b)$ for all $y\in\mathbb{R}$. If $f^{\prime}_{a,b}(\mathbf{K}^{s}_{u})=\gamma$ and $(a,b)\in{AB}_{c}$ then, by convexity of $f_{a,b}$, the tangent of $f$ at $y=\mathbf{K}^{s}_{u}$ satisfies for all $k\in\mathbb{N}$, $f(k)\geq f^{\prime}(\mathbf{K}^{s}_{u})(k-\mathbf{K}^{s}_{u})+f(\mathbf{K}^{s}_{u})$ which is the same as $ab^{k}+c\geq\gamma(k-\mathbf{K}^{s}_{u})+u_{\mathbf{K}^{s}_{u}}$. By definition of $\gamma$, we have for all $k\in(\mathbf{K}^{s}_{u},N_{c})$, $(\mathbf{K}^{s}_{u}-k)\gamma\leq u_{\mathbf{K}^{s}_{u}}-u_{k}$ and so $u_{k}\leq u_{\mathbf{K}^{s}_{u}}+(k-\mathbf{K}^{s}_{u})\gamma$. We conclude that if $f^{\prime}_{a,b}(\mathbf{K}^{s}_{u})=\gamma$ and $(a,b)\in{AB}_{c}$, $u_{k}\leq ab^{k}+c$ for all $k\in(\mathbf{K}^{s}_{u},N_{c}]$. Finally, we are looking for $(a,b)\in\mathbb{R}_{+}\times(0,1)$ such that :
$f_{a,b}(\mathbf{K}^{s}_{u})=u_{\mathbf{K}^{s}_{u}}\text{ i.e. }(a,b)\in{AB}_{c}\text{ and }f^{\prime}_{a,b}(\mathbf{K}^{s}_{u})=\gamma$
From $(a,b)\in{AB}_{c}$, we get $ab^{\mathbf{K}^{s}_{u}}=u_{\mathbf{K}^{s}_{u}}-c$. Hence, $f^{\prime}_{a,b}(\mathbf{K}^{s}_{u})=\gamma$ if and only if $\ln(b)=\frac{\gamma}{u_{\mathbf{K}^{s}_{u}}-c}<0$. So $b=\exp(\frac{\gamma}{u_{\mathbf{K}^{s}_{u}}-c})\in(0,1)$ and $a=(u_{\mathbf{K}^{s}_{u}}-c)\exp(-\frac{\mathbf{K}^{s}_{u}\gamma}{u_{\mathbf{K}^{s}_{u}}-c})>0$. ∎
In the proof of Theorem 1, we exhibit the parameters of the affine function. Unfortunately (but naturally) all the parameters depend on $\mathbf{K}^{s}_{u}$ and $u_{\mathbf{K}^{s}_{u}}$ what is exactly what we want to compute. Theorem 1 allows to construct an affine function in $\Gamma_{\rm f}(u)$ which is optimal. Optimal means that at the index $\mathbf{K}^{s}_{u}$, the upper-bound $ab^{\mathbf{K}^{s}_{u}}+c$ is equal to $u_{\mathbf{K}^{s}_{u}}$. Hence, we obtain $\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},x\mapsto ax+c,b)=\mathbf{K}^{s}_{u}$.
Corollary 1.
Let $u\in\Lambda$ such that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. Then, there exists a pair $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$ such that $u_{\mathbf{K}^{s}_{u}}=h_{\mathbf{K}^{s}_{u}}(\beta^{\mathbf{K}^{s}_{u}})$.
Theorem 1 proves that the nonemptiness of ${\mathrm{R}}_{u}^{s}$ is actually a sufficient condition for the existence of a usefully decreasing element in $\Gamma_{{\rm ev}}^{\rm dec}(u)$. Now, we also prove the necessity in Theorem 2. The necessity requires an auxiliary result formulated in Lemma 1.
Lemma 1.
Let $u\in\Lambda$. Let $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$ be usefully decreasing. Then:
  • the sequence $(\beta_{k}^{k})_{k\in\mathbb{N}}$ converges to 0 strictly decreasingly from $\mathfrak{m}(h,\beta)$;
  • for all $n\geq\mathfrak{m}(h,\beta)$, $\boldsymbol{\ell}(u)\leq h_{n}(0)$.
Proof.
Let us write $N=\mathfrak{m}(h,\beta)$. For all $n\geq N$, $10<\beta_{n+1}\leq\beta_{n}\leq\beta_{N}<$ 1 and thus $0<\beta_{n+1}^{n+1}<\beta_{n+1}^{n}\leq\beta_{n}^{n}\leq\beta_{N}^{n}$. Taking the limit over $n$ leads to $\boldsymbol{\ell}(\beta)=\lim_{n\to+\infty}\beta_{n}^{n}=0$.
[…]
Theorem 2.
Let $u\in\Lambda$. Then ${\mathrm{R}}_{u}^{s}\neq\emptyset$ if and only if there exists $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$ usefully decreasing.
Proof.
Theorem 1 asserts that ${\Gamma}_{0}^{\rm cst}(u)$ is nonempty when ${\mathrm{R}}_{u}^{s}$ is nonempty and therefore this implies the sufficiency. The necessity is a consequence of Lemma 1. Indeed, if there exists $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$ usefully decreasing, by definition, we can pick $n\in\mathcal{S}(u,h)\cap[\mathfrak{m}(h,\beta),+\infty)$. Therefore, $u_{n}>h_{n}(0)\geq\boldsymbol{\ell}(u)$ meaning that $n\in{\mathrm{R}}_{u}^{s}$. ∎
Finally, we can ask if we could pick non constant elements in $\Gamma(u)$ rather than the affine function furnished by Theorem 1. The answer is positive and the construction of the sequence is based on Lemma 2.
Lemma 2.
Let $u\in\Lambda$ such that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. Let us consider the triplet $(a,b,c)\in\mathbb{R}_{+}^{*}\times(0,1)\times\mathbb{R}$ constructed in Theorem. 1. Let us consider:
  1. 1.
    a decreasing sequence $\gamma:=(\gamma_{k})_{k\in\mathbb{N}}$ in $(c,\mathfrak{s}(u))$;
  2. 2.
    a decreasing sequence $\beta:=(\beta_{k})_{k\in\mathbb{N}}$ in $[b,1)$;
  3. 3.
    a decreasing sequence $\alpha:=(\alpha_{k})_{k\in\mathbb{N}}\in\Omega([0,1])^{\mathbb{N}}$ such that :
    $\inf_{k\in\mathbb{N}}\inf_{x\in(0,1]}\dfrac{\alpha_{k}(x)}{x}\geq a.$ (5)
Let us define for all $k\in\mathbb{N}$, $h_{k}:[0,1]\ni x\mapsto a_{k}(x)+\gamma_{k}$ and $h=(h_{k})_{k\in\mathbb{N}}$. Then, $(h,\beta)\in\Gamma_{0}^{\rm dec}(u)$ and $(h,\beta)$ is useful.
Proof.
We have from for all $k\in\mathbb{N}$, $\beta_{k}\in[b,1)$, $\alpha_{k}(\beta_{k}^{k})\geq a\beta_{k}^{k}\geq ab^{k}$. Thus, from Lemma 1, $h_{k}(\beta_{k}^{k})\geq ab^{k}+c\geq u_{k}$ and $(h,\beta)\in\Gamma(u)$. By assumptions on $\alpha$ and $\gamma$, $h$ is clearly decreasing and we conclude that $h\in{\Gamma}_{\rm f,{\rm ev}}^{\rm dec}(u)$. Finally, $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$. Moreover, since for all $k\in\mathbb{N}$ $h_{k}(0)=\gamma_{k}<\mathfrak{s}(u)=u_{\mathbf{K}^{s}_{u}}$, $\mathbf{K}^{s}_{u}\in\mathcal{S}(u,h)$ and $h$ is useful. ∎
Proposition 8.
Let $u\in\Lambda$ such that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. There exists $(h,\beta)\in\Gamma_{0}^{\rm dec}(u)$ useful such that $h$ and $\beta$ are not constant sequences.
Proof.
It suffices to apply Lemma 2 with for all $k\in\mathbb{N}$, $\alpha_{k}:x\mapsto x(k+1)^{-1}+a$, $\beta_{k}=\min\{b+(k+1)^{-1},(b+1)/2\}$ and $\gamma_{k}=\min\{c+(k+1)^{-1},(\mathfrak{s}(u)+c)/2\}$. ∎
Other possible sequences $\alpha_{k}$ in Lemma 2 can be chosen. Any sequence of functions of the form $\alpha_{k}:x\mapsto f(k)g(x)+a$ where $f$ is a decreasing strictly positive function and $g$ is a strictly increasing continuous positive function such that $g(x)\geq x^{t}$ with $t>1$.
3.3 Main properties of the upper boundfunctional $\mathfrak{F}_{u}$
Theorem 2 and Proposition 8 prove that we can find $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$ when ${\mathrm{R}}_{u}^{s}$ is nonempty. Hence, $\mathfrak{F}_{u}$ is finite somewhere when ${\mathrm{R}}_{u}^{s}$ is nonempty. Now, we present the main properties of the upper bound functional $(k,h,\beta)\to\mathfrak{F}_{u}(k,h,\beta)$: the increasing properties of $\mathfrak{F}_{u}$ with respect to $h$ and $\beta$ and the fact that $\mathfrak{F}_{u}(k,h,\beta)$ provides upper bounds of $\mathbf{K}^{s}_{u}$.
3.3.1 Increase properties of the upper bound functional $\mathfrak{F}_{u}$
The function $(h,\beta)\to\mathfrak{F}_{u}(\cdot,h,\beta)$ relies on the inverse of a function $h\in\Omega([0,1])$. Hence, to study increasing properties of $\mathfrak{F}_{u}(\cdot,h,\beta)$ with respect to $h$ we give some details on the decreasing property of the application $\Omega([0,1])\ni f\to f^{-1}$. First, we pay attention to the fact that we can only compare the inverses of two elements $f,g\in\Omega([0,1])$ on the values in the intersection of their range. Hence, we define, for all sets $J$ and all $\{h_{j},j\in J\}\subset\Omega([0,1])$:
$I_{[0,1]}\left(\{h_{j},j\in J\}\right)=\bigcap_{j\in J}[h_{j}(0),h_{j}(1)]=\left[\sup_{j\in J}h_{j}(0),\inf_{j\in J}h_{j}(1)\right]$ (6)
This property on the inverses is also extended to compute the inverses of minima and maxima of a finite family of elements of $\Omega([0,1])$. For a finite family $\{h_{i}:i\in I\}\subset\Omega([0,1])$, we write:
$\underline{h}_{I}:[0,1]\ni x\mapsto\min_{i\in I}h_{i}(x)\quad\text{ and }\quad\overline{h}_{I}:[0,1]\ni x\mapsto\max_{i\in I}h_{i}(x).$ (7)
and we define :
$\forall\,y\in[\underline{h}_{I}(0),\underline{h}_{I}(1)],\ \underline{I}(y):=\{i\in I:h_{i}(0)\leq y\}$ (8)
and
$\forall\,y\in[\overline{h}_{I}(0),\overline{h}_{I}(1)],\ \overline{I}(y):=\{i\in I:y\leq h_{i}(1)\}.$ (9)
The set $\underline{I}(y)$ excludes the functions $h_{i}$ for which $y$ is strictly less than $h_{i}(0)$ and so does not belong to $[h_{i}(0),h_{i}(1)]$. For those functions $h_{i}$, the inverse is not defined for this $y$ and we cannot use them to define the inverse of $\underline{h}_{I}$. The set $\overline{I}(y)$ has a similar use for $\overline{h}_{I}$.
Lemma 3.
Let $f,g\in\Omega([0,1])$ and $\{h_{i}:i\in I\}$ be a finite family in $\Omega([0,1])$. The following statements hold.
  1. 1.
    If $I_{[0,1]}\left(\{f,g\}\right)=\emptyset$, $f<g\iff f(0)<g(0)$.
  2. 2.
    If $I_{[0,1]}\left(\{f,g\}\right)\neq\emptyset$, $f\leq g\iff\forall\,y\in I_{[0,1]}\left(\{f,g\}\right),g^{-1}(y)\leq f^{-1}(y)$.
  3. 3.
    The functions $\underline{h}_{I},\overline{h}_{I}$ belong to $\Omega([0,1])$ and for all $y\in[\underline{h}_{I}(0),\underline{h}_{I}(1)]$ and all $z\in[\overline{h}_{I}(0),\overline{h}_{I}(1)]$:
    $\underline{h}_{I}^{-1}(y)=\max_{i\in\underline{I}(y)}h_{i}^{-1}(y)\quad\text{ and }\quad\overline{h}_{I}^{-1}(z)=\min_{i\in\overline{I}(z)}h_{i}^{-1}(z).$
Proof.
1. Suppose that $I_{[0,1]}\left(\{f,g\}\right)=\emptyset$. If $f(x)<g(x)$ for all $x\in[0,1]$, it obviously implies that $f(0)<g(0)$. Now suppose that $f(0)<g(0)$. As $I_{[0,1]}\left(\{f,g\}\right)=\emptyset$, we have for all $x\in[0,1]$ $f(x)\notin[g(0),g(1)]$. Thus either, $f(0)>g(1)$ or $f(1)<g(0)$. Then we exclude $f(0)>g(1)$ as $f(0)<g(0)$ and $g$ is strictly increasing on $[0,1]$. We conclude that $f(x)\leq f(1)<g(0)\leq g(x)$ for all $x\in[0,1]$ as $f$ and $g$ are strictly increasing on $[0,1]$.
2. Suppose that $I_{[0,1]}\left(\{f,g\}\right)\neq\emptyset$. Assume that $f\leq g$ on $[0,1]$. Let $y\in I_{[0,1]}\left(\{f,g\}\right)$. As $y\in[g(0),g(1)]$, $y=g(x)$ for some $x\in[0,1]$. Moreover, by assumption, we have $f(x)\leq g(x)$. As $y\in[f(0),f(1)]$, we can apply $f^{-1}$ which is strictly increasing to $g(x)=y$. We get $x\leq f^{-1}(g(x))$. As $x=g^{-1}(y)$, we conclude that $g^{-1}(y)\leq f^{-1}(y)$.
Now assume that for all $y\in I_{[0,1]}\left(\{f,g\}\right),g^{-1}(y)\leq f^{-1}(y)$. Let $x\in[0,1]$. We write $z=f(x)$ and thus $z\in[f(0),f(1)]$. Suppose that $z\in I_{[0,1]}\left(\{f,g\}\right)$, then $g^{-1}(z)\leq f^{-1}(z)$. Applying $g$, we get $z\leq g(f^{-1}(z))$ which is the same as $f(x)\leq g(x)$. Now suppose that $z\notin I_{[0,1]}\left(\{f,g\}\right)$ meaning that $z\notin[g(0),g(1)]$. If $z<g(0)$, then we have $f(x)<g(0)\leq g(x)$ as $g$ increases on $[0,1]$. Assume now that $z=f(x)>g(1)$. As $I_{[0,1]}\left(\{f,g\}\right)$ is nonempty, there exists $u\in[0,x)$ such that $f(u)\leq g(1)$. By continuity of $f$, there exists $v\in[u,x)$ such that $f(v)=g(1)$. As $f(v)\in I_{[0,1]}\left(\{f,g\}\right)$, we have $g^{-1}(f(v))\leq f^{-1}(f(v))=v$ and thus $g^{-1}(g(1))=1\leq v<x\leq 1$. This implies that $z$ cannot greater than $g(1)$. This ends the proof.
3. The minimum and the maximum of a finite family of strictly increasing continuous functions is still strictly increasing and continuous; hence $\underline{h}_{I},\overline{h}_{I}\in\Omega([0,1])$. Let $y\in[\underline{h}_{I}(0),\underline{h}_{I}(1)]$. Then there exists $x\in[0,1]$ such that $y=\underline{h}_{I}(x)$. As $I$ is finite, there exists $\ell\in I$ such that $\underline{h}_{I}(x)=h_{\ell}(x)$. The index $\ell$ belongs to $\underline{I}(y)$ as for all $i\notin\underline{I}(y)$, $y<h_{i}(0)\leq h_{i}(x)$ for all $x\in[0,1]$. Now let $i\in\underline{I}(y)$. Then $y\in[h_{i}(0),\underline{h}_{I}(1)]\subseteq I_{[0,1]}\left(\{h_{i},\underline{h}_{I}\}\right)$. From the second statement, $h_{i}^{-1}(y)\leq\underline{h}_{I}^{-1}(y)$. Finally, $\max_{i\in\underline{I}(y)}h_{i}^{-1}(y)\leq\underline{h}_{I}^{-1}(y)=x=h_{\ell}^{-1}(y)\leq\max_{i\in\underline{I}(y)}h_{i}^{-1}(y)$ and we conclude that $\underline{h}_{I}^{-1}(y)=\max_{i\in\underline{I}(y)}h_{i}^{-1}(y)$. The proof for $\overline{h}_{I}^{-1}$ is similar. ∎
The set $\Omega([0,1])^{\mathbb{N}}$ is naturally endowed with standard lattice operations of sequence spaces. We recall that for $g,h\in\Omega([0,1])^{\mathbb{N}}$, $g\leq h$ iff $g_{n}(x)\leq h_{n}(x)$ for all $n\in\mathbb{N}$ for all $x\in[0,1]$. Hence, using the notation of Equation (7), we define the minimum and the maximum of a finite family $\{h_{(i)},i\in I\}\subset\Omega([0,1])^{\mathbb{N}}$ as follows:
$\begin{array}[]{lc}&\underline{h}_{(I)}:=\min_{i\in I}h_{(i)}=\left([0,1]\ni x\to\min_{i\in I}h_{(i),k}(x)\right)_{k\in\mathbb{N}}=\left(\underline{h}_{(I),k}\right)_{k\in\mathbb{N}}\\ \text{ and }&\overline{h}_{(I)}:=\max_{i\in I}h_{(i)}=\left([0,1]\ni x\to\max_{i\in I}h_{(i),k}(x)\right)_{k\in\mathbb{N}}=\left(\overline{h}_{(I),k}\right)_{k\in\mathbb{N}}.\end{array}$
Moreover, for a finite family $\{\beta_{(i)}:i\in I\}\subset(0,1)^{\mathbb{N}}$, we denote by:
$\underline{\beta}_{(I)}:=\left(\min_{i\in I}\beta_{(i),k}\right)_{k\in\mathbb{N}}\text{ and }\overline{\beta}_{(I)}:=\left(\max_{i\in I}\beta_{(i),k}\right)_{k\in\mathbb{N}}$
respectively the minimum and the maximum of $\{\beta_{(i)}:i\in I\}$.
It can be noticed that any element of $\Omega([0,1])^{\mathbb{N}}$ greater than an element of $\Gamma_{\rm f}(u)$ belongs to $\Gamma_{\rm f}(u)$. Besides, the set function $\Gamma_{\rm f}(u)\ni h\to\Gamma_{\rm s}(u,h)$ is order-preserving with respect to the order defined on $\Omega([0,1])^{\mathbb{N}}$. Moreover, as the element of $\Omega([0,1])$ are increasing functions, any element of $(0,1)^{\mathbb{N}}$ greater than an element of $\Gamma_{\rm s}(u,h)$ belongs to $\Gamma_{\rm s}(u,h)$. Those simple facts are presented in Lemma 4.
Lemma 4.
Let $u\in\mathbb{R}^{\mathbb{N}}$ and $h\in\Gamma_{\rm f}(u)$.
  1. 1.
    Let $h\in\Omega([0,1])^{\mathbb{N}}$ such that $g\leq h$, then $h\in\Gamma_{\rm f}(u)$ and $\Gamma_{\rm s}(u,g)\subseteq\Gamma_{\rm s}(u,h)$.
  2. 2.
    Let $\beta\in\Gamma_{\rm s}(u,h)$. Then all $\beta^{\prime}\in(0,1)^{\mathbb{N}}$ such that $\beta\leq\beta^{\prime}$ belong to $\Gamma_{\rm s}(u,h)$.
Remark 3.
We should mention that the two statements in Lemma 4 refer to a well-known notion from ordered sets theory. More precisely, $\Gamma_{\rm f}(u)$ and $\Gamma_{\rm s}(u,h)$ are said to be upward closed (or upset or order filter). The interested reader can consult [DP90] or [Rom08].
Proposition 9.
Let $u\in\mathbb{R}^{\mathbb{N}}$. The following assertions hold:
  • For all finite sets $I$, $\left\{(h_{(i)},\beta_{(i)}):i\in I\right\}\subset\Gamma(u)$:
    $\overline{\beta}_{(I)}\in\bigcap_{i\in I}\Gamma_{\rm s}(u,h_{(i)}).$
  • Let $I$ be finite and $\{h_{(i)}:i\in I\}\subset\Gamma_{\rm f}(u)$. Then $\underline{h}_{(I)}$ and $\overline{h}_{(I)}$ belong to $\Gamma_{\rm f}(u)$ and
    $\Gamma_{\rm s}\left(u,\underline{h}_{(I)}\right)=\bigcap_{i\in I}\Gamma_{\rm s}\left(u,h_{(i)}\right)\subseteq\Gamma_{\rm s}\left(u,\overline{h}_{(I)}\right).$
    If, moreover, for all $i\in I$, $h_{(i)}\in{\Gamma}_{\rm f,{\rm ev}}^{\rm dec}(u)$, then $\underline{h}_{(I)}$ and $\overline{h}_{(I)}$ belong to ${\Gamma}_{\rm f,{\rm ev}}^{\rm dec}(u)$ and $\mathfrak{m}(\underline{h}_{(I)})$ and $\mathfrak{m}(\overline{h}_{(I)})$ are smaller than ${\max_{i\in I}\mathfrak{m}(h_{(i)})}$.
Proof.
The first statement is a direct consequence of the first assertion of Lemma 4.
Let us prove the second statement. First, $\underline{h}_{(I)}$ and $\overline{h}_{(I)}$ belong $\Omega([0,1])^{\mathbb{N}}$ from the third statement of Lemma 3. The fact that $\overline{h}_{(I)}$ belongs to $\Gamma_{\rm f}(u)$ is a consequence of the first assertion of Lemma 4. Now as for all $i\in I$, $h_{(i)}\in\Gamma_{\rm f}(u)$, we have for all $i\in I$, for all $k\in\mathbb{N}$, $u_{k}\leq h_{(i),k}(\beta_{k}^{k})$ for all $\beta\in\cap_{i\in I}\Gamma_{\rm s}\left(u,h_{(i)}\right)$. Hence, for all $k\in\mathbb{N}$, $u_{k}\leq\underline{h}_{(I),k}(\beta_{k}^{k})$ for all $\beta\in\cap_{i\in I}\Gamma_{\rm s}(u,h_{(i)})$. This proves that $\underline{h}_{(I)}\in\Gamma_{\rm f}(u)$ and $\cap_{i\in I}\Gamma_{\rm s}(u,h_{(i)})\subseteq\Gamma_{\rm s}(u,\underline{h}_{(I)})$. The converse inclusion is a consequence of the first statement of Lemma 4. Now suppose that $h_{(i)}\in{\Gamma}_{\rm f,{\rm ev}}^{\rm dec}(u)$ for all $i\in I$. Then, we have, for all $i\in I$, for all $x\in[0,1]$, $h_{(i),k}(x)\leq h_{(i),k+1}(x)$ for all $k\geq\mathfrak{m}(h_{(i)})$. This implies that for all $x\in[0,1]$, $\underline{h}_{(I),k}(x)\leq\underline{h}_{(I),k+1}(x)$ for all $k\geq\max_{i\in I}\mathfrak{m}(h_{(i)})$ and $\underline{h}_{(I)}\in{\Gamma}_{\rm f,{\rm ev}}^{\rm dec}(u)$ with $\max_{i\in I}\mathfrak{m}(h_{(i)})\geq\mathfrak{m}(\underline{h}_{(I)})$. ∎
Lemma 4 and Proposition 9 show the increase properties of $\Gamma_{\rm s}(u,h)$ with respect to $h$. From those results, we can derive increase properties of $\mathfrak{F}_{u}$ with respect to $h$ and $\beta$.
Proposition 10.
Let $u\in\mathbb{R}^{\mathbb{N}}$. The following properties hold:
  1. 1.
    Let $h\in\Gamma_{\rm f}(u)$ be useful and $k\in\mathcal{S}(u,h)$. Let $\left\{\beta_{(i)},i\in I\right\}$ a finite family in $\Gamma_{\rm s}(u,h)$. Then
    $\mathfrak{F}_{u}\left(k,h,\underline{\beta}_{(I)}\right)\leq\min_{i\in I}\mathfrak{F}_{u}\left(k,h,\beta_{(i)}\right).$
  2. 2.
    Let $\left\{h_{(i)}:i\in I\right\}$ be a finite family in $\Gamma_{\rm f}(u)$ such that at least one element is useful. Let $k\in\cup_{i\in I}\mathcal{S}\left(u,h_{(i)}\right)$. Let $\beta\in\Gamma_{\rm s}\left(u,\underline{h}_{(I)}\right)$. Then
    $\mathfrak{F}_{u}\left(k,\underline{h}_{(I)},\beta\right)\leq\min_{i\in I}\mathfrak{F}_{u}\left(k,h_{(i)},\beta\right).$
Proof.
1. The result follows from the fact that the function $(0,1)\ni s\mapsto\ln(h_{k}^{-1}(u_{k}))/\ln(s)$ is increasing as $\ln(h_{k}^{-1}(u_{k}))$ is negative.
2. First from Proposition 9, $\underline{h}_{I}\in\Gamma_{\rm f}(u)$ and $\beta\in\cap_{i\in I}\Gamma_{\rm s}(u,h_{(i)})$. Second, from Lemma 3, for all $k\in\mathbb{N}$, ${\underline{h}_{(I)}}_{k}^{-1}(u_{k})=\max_{i\in\underline{I}_{k}(u_{k})}h_{(i),k}^{-1}(u_{k})$ where $\underline{I}_{k}(u_{k})=\{i\in I:h_{(i),k}(0)\leq u_{k}\}$. Note that $\underline{I}_{k}(u_{k})$ is nonempty as it contains all $i\in I$ such that $k\in\mathcal{S}\left(u,h_{(i)}\right)$. Moreover, ${\underline{h}}_{(I),k}^{-1}(u_{k})\geq h_{(j),k}^{-1}(u_{k})$ for all $j\in I$ such that $h_{(j),k}(0)<u_{k}$ and so ${\underline{h}}_{(I),k}^{-1}(u_{k})>0$. Let $i\in I$. If $k\notin\mathcal{S}(u,h_{(i)})$, $\mathfrak{F}_{u}(k,h_{(i)},\beta)=+\infty$. Now assume that $k\in\mathcal{S}(u,h_{(i)})$, $h_{(i),k}^{-1}(u_{k})\leq{\underline{h}}_{(I),k}^{-1}(u_{k})$. Now applying $\ln$ and dividing by $\ln(\beta_{k})<0$, we get the desired result. ∎
Corollary 2.
Let $u\in\mathbb{R}^{\mathbb{N}}$. Let $h\in\Gamma(u)$. Then, for all useful $g\in\Gamma(u)$ such that $g\leq h$, we have for all $k\in\mathcal{S}(u,g)$ and all $\beta\in\Gamma_{\rm s}(u,g)$:
$\mathfrak{F}_{u}(k,g,\beta)\leq\mathfrak{F}_{u}(k,h,\beta).$
The first statement of Proposition 10 has to be read as follows : when a sequence $h\in\Gamma_{\rm f}(u)$ has been found, we have to choose the smallest possible $\beta\in\Gamma_{\rm s}(u,h)$ to obtain the smallest possible value of $\mathfrak{F}_{u}(k,h,\beta)$. The second statement has a similar meaning in a functional sense: if we have at hand several functions in $\Gamma_{\rm f}(u)$ sharing a same sequence $\beta$, we have to consider the minimum of these functions to obtain the smallest possible value of $\mathfrak{F}_{u}(k,h,\beta)$.
Remark 4.
In the proof of Proposition 5, the pair $(h,\beta)\in\Gamma_{0}^{\rm dec}(u)$ constructed from $(g,\gamma)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$ is greater than $(g,\gamma)$. In this case, as proved in Proposition 10, to use $(h,\beta)$ instead of $(g,\gamma)$ makes increase the value of $\mathfrak{F}_{u}(k,\cdot,\cdot)$ i.e. $\mathfrak{F}_{u}(k,g,\gamma)\leq\mathfrak{F}_{u}(k,h,\beta)$ for all $k\in\mathcal{S}(u,g)\cap[0,\mathfrak{m}(g,\gamma)]$.
3.3.2 The upper bound functional $\mathfrak{F}_{u}$ as an upper bound of $\mathbf{K}^{s}_{u}$
Here, we prove that using eventually and usefully decreasing pairs $(h,\beta)$ permits to prove that $\mathfrak{F}_{u}(k,h,\beta)$ is an upper bound of $\mathbf{K}^{s}_{u}$ for all $k\geq\mathfrak{m}(h,\beta)$. First, we prove some useful properties satisfied by $\mathfrak{F}_{u}$.
Theorem 3.
Let $u\in\mathbb{R}^{\mathbb{N}}$ and $(h,\beta)\in\Gamma(u)$ useful. Let $k\in\mathcal{S}(u,h)$. The following statements hold.
  1. 1.
    $\mathfrak{F}_{u}(k,h,\beta)$ is well-defined and, moreover, strictly positive if $k\neq 0$ and $\mathfrak{F}_{u}(0,h,\beta)=0$ if and only if $u_{0}=h_{0}(1)$.
  2. 2.
    The equalities hold:
    $\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1=\min\{j\in\mathbb{N}:h_{k}(\beta_{k}^{j})<u_{k}\}=\min\{j\in\mathbb{N}:\beta_{k}^{j}<h_{k}^{-1}(u_{k})\}$
  3. 3.
    $\mathfrak{F}_{u}(k,h,\beta)\geq k$;
  4. 4.
    Assume that $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$ and $k\geq\mathfrak{m}(h,\beta)$. For all $j>\mathfrak{F}_{u}(k,h,\beta)$, $u_{j}<u_{k}$;
  5. 5.
    Assume that $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$ and $k\geq\mathfrak{m}(h,\beta)$. Let $j\in[\mathfrak{m}(h,\beta),k]\cap\mathbb{N}$. If $u_{j}\leq u_{k}$ then $\mathfrak{F}_{u}(k,h,\beta)\leq\mathfrak{F}_{u}(j,h,\beta)$. If moreover, $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm cst}(u)$ and $j,k\geq\mathfrak{c}(h,\beta)$, $u_{j}\leq u_{k}\implies\mathfrak{F}_{u}(k,h,\beta)\leq\mathfrak{F}_{u}(j,h,\beta)$ (the condition $j\leq k$ can be dropped).
Proof.
1. Let $(h,\beta)\in\Gamma(u)$. As $h\in\Gamma_{\rm f}(u)$ and we only consider $u_{k}>h_{k}(0)$, we have $h_{k}^{-1}(u_{k})>0$. We also have, if $k\in\mathbb{N}^{*}$, $u_{k}<h_{k}(1)$ meaning that $h_{k}^{-1}(u_{k})\in(0,1)$. If $u_{0}<h_{0}(1)$, the strict inequality $u_{k}<h_{k}(1)$ is also valid at $k=0$. As for all $k\in\mathbb{N}$, $\beta_{k}\in(0,1)$, $\ln(h_{k}^{-1}(u_{k}))$ and $\ln(\beta_{k})$ are both strictly negative and then the result holds. It is easy to see that $u_{0}=h_{0}(1)$ iff $\mathfrak{F}_{u}(0,h,\beta)=0$.
2. As $(h,\beta)\in\Gamma(u)$ and $k\in\mathcal{S}(u,h)$, there exists $j\in\mathbb{N}$ such that $h_{k}(\beta_{k}^{j})<u_{k}$. Then the sets $\{j\in\mathbb{N}:h_{k}(\beta_{k}^{j})<u_{k}\}$ and $\{j\in\mathbb{N}:\beta_{k}^{j}<h_{k}^{-1}(u_{k})\}$ are equal and nonempty. Let $j\in\mathbb{N}$ s.t. $\beta_{k}^{j}<h_{k}^{-1}(u_{k})$. Applying the natural logarithm and dividing by $\ln(\beta_{k})<0$ leads to $j>\mathfrak{F}_{u}(k,h,\beta)$. As $j\in\mathbb{N}$, we have $j\geq\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1$. Since $\beta_{k}\in(0,1)$, $\ln(\beta_{k})(\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1)<\mathfrak{F}_{u}(k,h,\beta)\ln(\beta_{k})=\ln(h_{k}^{-1}(u_{k}))$ and $\beta_{k}^{\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1}<h_{k}^{-1}(u_{k})$. Hence $\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1$ belongs to $\{j\in\mathbb{N}:\beta_{k}^{j}<h_{k}^{-1}(u_{k})\}$ and is its minimum.
3. As $(h,\beta)\in\Gamma(u)$ and $k\in\mathcal{S}(u,h)$, we have $10<h_{k}^{-1}(u_{k})\leq\beta_{k}^{k}<$ 1. Applying the natural logarithm and dividing by $\ln(\beta_{k})<0$ leads to the result.
4. Let $j>\mathfrak{F}_{u}(k,h,\beta)$ then $j\geq\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1>k$ and, as $h_{k}$ is increasing, $h_{k}(\beta_{k}^{j})\leq h_{k}(\beta_{k}^{\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1})<u_{k}$ from the second statement. Moreover, by definition, as $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$ and $j>k\geq\mathfrak{m}(h,\beta)$, we have $h_{j}(x)\leq h_{k}(x)$ for all $x\in[0,1]$ and $10<\beta_{j}\leq\beta_{k}<$ 1. Then, $h_{j}(\beta_{j}^{j})\leq h_{k}(\beta_{j}^{j})$ and as $h_{k}$ is increasing $h_{k}(\beta_{j}^{j})\leq h_{k}(\beta_{k}^{j})$. It follows that $u_{j}\leq h_{j}(\beta_{j}^{j})<u_{k}$.
5. Let $j\geq[\mathfrak{m}(h,\beta),k]\cap\mathbb{N}$ and assume that $u_{j}\leq u_{k}$. If $j\notin\mathcal{S}(u,h)$, the result holds as $\mathfrak{F}_{u}(j,h,\beta)=+\infty$. Now suppose that $j\in\mathcal{S}(u,h)$. As $k\leq j\geq\mathfrak{m}(h,\beta)$, we have $h_{k}(x)\leq h_{j}(x)$ for all $x\in[0,1]$. This implies that $h_{k}(0)\leq h_{j}(0)$ and $h_{k}(1)\leq h_{j}(1)$. As $j\in\mathcal{S}(u,h)$, we have $h_{j}(0)<u_{j}$ and as $h\in\Gamma_{\rm f}(u)$, we have $u_{k}<h_{k}(1)$. By assumption, $u_{j}\leq u_{k}$ then $u_{j},u_{k}\in I_{[0,1]}\left(\{h_{k},h_{j}\}\right)=[h_{j}(0),h_{k}(1)]$. By Lemma 3, we have $h_{j}^{-1}(u_{j})\leq h_{k}^{-1}(u_{j})$. Hence, as $\ln$ and $h_{k}^{-1}$ are increasing, $\ln(h_{j}^{-1}(u_{j}))\leq\ln(h_{k}^{-1}(u_{j}))\leq\ln(h_{k}^{-1}(u_{k}))$. Moreover, as $\beta$ is decreasing from $\mathfrak{m}(h,\beta)$, we get $\ln(\beta_{j})/\ln(\beta_{k})\in(0,1)$. Dividing by $\ln(\beta_{j})<0$ we have $\mathfrak{F}_{u}(j,h,\beta)\geq\ln(h_{k}^{-1}(u_{k})/\ln(\beta_{j})\geq(\ln(h_{k}^{-1}(u_{k})/\ln(\beta_{j}))\times(\ln(\beta_{j})/\ln(\beta_{k}))=\mathfrak{F}_{u}(k,h,\beta)$.
Assume that there exists $(g,\gamma)\in\Omega([0,1])\times(0,1)$ such that $(h_{n},\beta_{n})=(g,\gamma)$ for all $n\geq\mathfrak{c}(h,\beta)$. The conclusion follows from the fact that $[g(0),g(1)]\ni x\to(\ln(g^{-1}(x)))/\ln(\gamma)$ is decreasing as $\gamma\in(0,1)$. ∎
Remark 5.
The second statement of Theorem 3 may suggest that $\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1$ is an extension of stopping times on positive integers [Ter76] or first passage time [Tao22] that appear in the study of Syracuse sequences. Indeed, $\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1$ is the smallest integer $j$ such that $h_{k}(\beta_{k}^{j})<u_{k}$ whereas $u_{k}\leq h_{k}(\beta_{k}^{k})$ as $(h,\beta)\in\Gamma(u)$. When $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$, this is confirmed by the fourth statement for which we know that after $\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor+1$, the terms of $u$ are strictly smaller than $u_{k}$. Without the auxiliary sequence characterized by $(h,\beta)$, it is not possible to establish such results.
Corollary 3.
Let $u\in\Lambda$ such that ${\mathrm{R}}_{u}^{s}\neq\emptyset$. Let $(h,\beta)\in\Gamma_{0}^{\rm dec}(u)$ be useful. Then:
  1. 1.
    For all $j\in\mathbb{N}$ such that $j>\mathfrak{F}_{u}(k,h,\beta)$, $u_{j}<u_{k}$;
  2. 2.
    Let $j\in\mathcal{S}(u,h)$. If $j\leq k$ and $u_{j}\leq u_{k}$ then $\mathfrak{F}_{u}(k,h,\beta)\leq\mathfrak{F}_{u}(j,h,\beta)$. If moreover, $(h,\beta)\in{\Gamma}_{0}^{\rm cst}(u)$, $u_{j}\leq u_{k}\implies\mathfrak{F}_{u}(k,h,\beta)\leq\mathfrak{F}_{u}(j,h,\beta)$.
Remark 6.
We insist on the fact that if $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$ and not in $\Gamma_{{\rm ev}}^{\rm cst}(u)$, we cannot guarantee that $\mathfrak{F}_{u}(k,h,\beta)\leq\mathfrak{F}_{u}(j,h,\beta)$ when $u_{j}\leq u_{k}$ and $j>k$. This means, when we compute $\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)$, we may find some $j>\mathbf{K}^{s}_{u}$ such that $\mathfrak{F}_{u}(j,h,\beta)$ is strictly smaller than $\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)$.
Proposition 11.
Let $u\in\mathbb{R}^{\mathbb{N}}$ and $(h,\beta)\in\Gamma(u)$. The following statements hold:
  1. 1.
    For all $j\geq\mathbf{K}^{s}_{u}$, $\mathbf{K}^{s}_{u}\leq\mathfrak{F}_{u}(j,h,\beta)$;
  2. 2.
    Assume that $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$ and is usefully decreasing, then:
    1. (a)
      for all $j\geq\mathfrak{m}(h,\beta)$, we have $\mathbf{K}^{s}_{u}\leq\mathfrak{F}_{u}(j,h,\beta)$;
    2. (b)
      if $\mathbf{K}^{s}_{u}\geq\mathfrak{m}(h,\beta)$, we have for all $j\in[\mathfrak{m}(h,\beta),\mathbf{K}^{s}_{u}]$:
      $\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)\leq\mathfrak{F}_{u}(j,h,\beta).$
  3. 3.
    If, moreover $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm cst}(u)$ (and is still usefully decreasing) and $\mathbf{K}^{s}_{u}\geq\mathfrak{c}(h,\beta)$, then we have:
    $\min_{k\in\mathcal{S}(u,h)\cap[\mathfrak{m}(h,\beta),+\infty)}\mathfrak{F}_{u}(k,h,\beta)=\mathfrak{F}_{u}(\mathbf{K}^{s}_{\nu},h,\beta)\enspace.$
Proof.
1. Let $j\geq\mathbf{K}^{s}_{u}$. If $j\notin\mathcal{S}(u,h)$, then $\mathfrak{F}_{u}(j,h,\beta)=+\infty$ and the inequality holds. Now, assume that $(h,\beta)$ is useful and $j\in\mathcal{S}(u,h)$. From the third statement of Theorem 3, $\mathfrak{F}_{u}(j,h,\beta)\geq j\geq\mathbf{K}^{s}_{u}$.
2. (a) Let assume that $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$. Let $j\geq\mathfrak{m}(h,\beta)$. If $j\notin\mathcal{S}(u,h)$, the inequality trivially holds. Now, assume that $j\in\mathcal{S}(u,h)$. If $\mathbf{K}^{s}_{u}>\mathfrak{F}_{u}(j,h,\beta)$, from the fourth statement of Theorem 3, $u_{\mathbf{K}^{s}_{u}}<u_{j}$. This contradicts that $u_{\mathbf{K}^{s}_{u}}=\mathfrak{s}(u)$.
(b) Let us suppose that $\mathbf{K}^{s}_{u}\geq\mathfrak{m}(h,\beta)$. Let $j\in[\mathfrak{m}(h,\beta),\mathbf{K}^{s}_{u}]$. If $j\notin\mathcal{S}(u,h)$, the inequality holds. Suppose that $j\in\mathcal{S}(u,h)$. Then, as $u_{j}\leq u_{\mathbf{K}^{s}_{u}}=\mathfrak{s}(u)$ and $j\leq\mathbf{K}^{s}_{u}$, from the last statement of Theorem 3, we have $\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)\leq\mathfrak{F}_{u}(j,h,\beta)$.
3. First, as $\Gamma_{{\rm ev}}^{\rm cst}(u)\subseteq\Gamma_{{\rm ev}}^{\rm dec}(u)$, $\mathfrak{m}(h,\beta)$ is defined for all $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm cst}(u)$. As $(h,\beta)$ can be decreasing before being constant, we have $\mathfrak{m}(h,\beta)\leq\mathfrak{c}(h,\beta)$. Let $(g,\gamma)\in\Omega([0,1])\times(0,1)$ such that $(h_{n},\beta_{n})=(g,\gamma)$ for all $n\geq\mathfrak{c}(h,\beta)$. We prove that $\mathbf{K}^{s}_{u}\in\mathcal{S}(u,h)$. As $(h,\beta)$ is usefully decreasing, $\mathcal{S}(u,h)\cap[\mathfrak{m}(h,\beta),+\infty)$ is nonempty. If there exists $j\in\mathcal{S}(u,h)\cap[\mathfrak{m}(h,\beta),+\infty)$ such that $j<\mathbf{K}^{s}_{u}$, then, as $j\geq\mathfrak{m}(h,\beta)$ and $\mathbf{K}^{s}_{u}\in\operatorname{Argmax}(u)$, we have $u_{\mathbf{K}^{s}_{u}}\geq u_{j}>h_{j}(0)\geq h_{\mathbf{K}^{s}_{u}}(0)$. If there exists $j\in\mathcal{S}(u,h)\cap[\mathfrak{m}(h,\beta),+\infty)$ such that $j>\mathbf{K}^{s}_{u}$, then, as $j>\mathbf{K}^{s}_{u}\geq\mathfrak{c}(h,\beta)$ and $\mathbf{K}^{s}_{u}\in\operatorname{Argmax}(u)$, we have $u_{\mathbf{K}^{s}_{u}}\geq u_{j}>h_{j}(0)=g(0)=h_{\mathbf{K}^{s}_{u}}(0)$.
Now let $k\in\mathcal{S}(u,h)\cap[\mathfrak{m}(h,\beta),+\infty)$. If $k\leq\mathbf{K}^{s}_{u}$, then as $\mathbf{K}^{s}_{u}\geq\mathfrak{m}(h,\beta)$, the assertion (b) of the second statement can be used and we have $\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)\leq\mathfrak{F}_{u}(k,h,\beta)$. Now suppose that $k>\mathbf{K}^{s}_{u}$. We can use the fifth statement of Theorem 3 as $k>\mathbf{K}^{s}_{u}\geq\mathfrak{c}(h,\beta)$, $u_{k}<u_{\mathbf{K}^{s}_{u}}$, $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm cst}(u)$ and $\mathbf{K}^{s}_{u}\in\mathcal{S}(u,h)$.
Corollary 4.
Let $u\in\Lambda$ with ${\mathrm{R}}_{u}^{s}\neq\emptyset$ and $(h,\beta)\in\Gamma_{0}^{\rm dec}(u)$ useful. The following statements hold:
  1. 1.
    For all $j\in\mathbb{N}$, $\mathbf{K}^{s}_{u}\leq\mathfrak{F}_{u}(j,h,\beta)$.
  2. 2.
    For all $j\leq\mathbf{K}^{s}_{u}$, $\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)\leq\mathfrak{F}_{u}(j,h,\beta)$.
  3. 3.
    If, moreover $(h,\beta)\in{\Gamma}_{0}^{\rm cst}(u)$, then we have:
    $\min_{k\in\mathcal{S}(u,h)}\mathfrak{F}_{u}(k,h,\beta)=\mathfrak{F}_{u}(\mathbf{K}^{s}_{\nu},h,\beta)\enspace.$
The third statement of of Corollary 4 can be directly used in a very particular context. Suppose that we have an explicit closed formula for $u_{k}$ i.e. there exists a function $f:\mathbb{R}_{+}\mapsto\mathbb{R}$ for all $k\in\mathbb{N}$ such that $u_{k}=f(k)$. Suppose that we have $(h,\beta)\in{\Gamma}_{0}^{\rm cst}(u)$ and so it is identified as an element of $\Omega([0,1])\times(0,1)$. The function $h^{-1}$ could be used to regularize and strictly convexify meaning that $(\ln(\beta))^{-1}\times\ln\circ h^{-1}\circ f$ is smooth and strictly convex. We thus obtain with some calculus the exact value of $\mathbf{K}^{s}_{\nu}$ as the integer part (or integer part +1) of the zero of the derivative of $\mathbb{R}_{+}\ni s\to\ln(h^{-1}(f(s)))/(\ln(\beta))$.
We terminate the presentation of the main properties of $\mathfrak{F}_{u}$ by a complete characterization of $\mathbf{K}^{s}_{u}$ from the upper bound function $\mathfrak{F}_{u}$.
Theorem 4 (Formula Optimality).
For all $u\in\mathbb{R}^{\mathbb{N}}$, we have the following formula:
$\mathbf{K}^{s}_{u}=\inf_{(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)}\ \inf_{\begin{subarray}{c}k\in\mathcal{S}(u,h)\\ k\geq\mathfrak{m}(h,\beta)\end{subarray}}\mathfrak{F}_{u}(k,h,\beta)=\inf_{(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)}\ \inf_{\begin{subarray}{c}k\in\mathcal{S}(u,h)\\ k\geq\mathfrak{m}(h,\beta)\end{subarray}}\dfrac{\ln(h^{-1}(u_{k}))}{\ln(\beta)}$
and the $\inf$ can be replaced by $\min$ when ${\mathrm{R}}_{u}^{s}\neq\emptyset$.
Proof.
Let us define
$F:=\inf_{(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)}\ \inf_{\begin{subarray}{c}k\in\mathcal{S}(u,h)\\ k\geq\mathfrak{m}(h,\beta)\end{subarray}}\mathfrak{F}_{u}(k,h,\beta).$
If $u\notin\Lambda$, then from Proposition 7, $\Gamma_{{\rm ev}}^{\rm dec}(u)=\emptyset$ and thus $F=+\infty$. Moreover, from Proposition 2, $\mathbf{K}^{s}_{u}=+\infty$.
Suppose that $u\in\Lambda$. From Proposition 7, $\Gamma_{{\rm ev}}^{\rm dec}(u)\neq\emptyset$. Assume that $\Delta^{s}_{u}=\emptyset$. This is the same as $\mathbf{K}^{s}_{u}=+\infty$. From Proposition 3, ${\mathrm{R}}_{u}^{s}=\emptyset$ and from Theorem 2, there does not exist a usefully decreasing element i.e. for all $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$, $\mathcal{S}(u,h)\cap[\mathfrak{m}(h,\beta),+\infty)=\emptyset$ which leads to $F=+\infty$. Assume that $\Delta^{s}_{u}\neq\emptyset$ and thus ${\mathrm{R}}_{u}^{s}\neq\emptyset$. From the second statement of Proposition 11, $\mathbf{K}^{s}_{u}\leq F$. Moreover, from Corollary 1, there exists $(h,\beta)\in\Gamma_{0}^{\rm dec}(u)$ such that $u_{\mathbf{K}^{s}_{u}}=h_{\mathbf{K}^{s}_{u}}(\beta_{\mathbf{K}^{s}_{u}}^{\mathbf{K}^{s}_{u}})>h_{\mathbf{K}^{s}_{u}}(0)$ and $F<+\infty$. Taking the inverse of $h_{\mathbf{K}^{s}_{u}}$ on $[0,1]$ and the natural logarithm, we get $\mathbf{K}^{s}_{u}=\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)$. By definition of $F$, $F\leq\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)=\mathbf{K}^{s}_{u}$. We conclude that $\mathbf{K}^{s}_{u}=F$. ∎
3.4 Resolution of Problem 1 using the upper bound functional $\mathfrak{F}_{u}$
As the upper bound function $\mathfrak{F}_{u}$ provides an upper bound of $\mathbf{K}^{s}_{u}$ for a given $u\in\Lambda$ with ${\mathrm{R}}_{u}^{s}\neq\emptyset$, we can solve Problem 1 in finite time truncating $u$ at $\lfloor\cdot\rfloor\circ\mathfrak{F}_{u}$. This is formally described in Theorem 5. The definition of the truncation index depends on the quality of the upper bound over $u$.
Theorem 5 (Getting a solution of Problem 1).
Let $u\in\Lambda$ with ${\mathrm{R}}_{u}^{s}\neq\emptyset$. Let $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$ be useful decreasing for $u$. Then :
$\max_{n\in\mathbb{N}}u_{n}=\max\{u_{n}:n=0,\ldots,\mathfrak{m}(h,\beta),\ldots,\min_{k\in\mathcal{S}(u,h)\cap[\mathfrak{m}(h,\beta),+\infty)}\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor\}.$
If moreover, $(h,\beta)\in\Gamma_{0}^{\rm dec}(u)$, then:
$\max_{n\in\mathbb{N}}u_{n}=\max\{u_{n}:n=0,\ldots,\min_{k\in\mathcal{S}(u,h)}\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor\}.$
If, finally, $(h,\beta)\in{\Gamma}_{0}^{\rm cst}(u)$, then:
$\max_{n\in\mathbb{N}}u_{n}=\max\{u_{n}:n=0,\ldots,\lfloor\mathfrak{F}_{u}(\mathbf{K}^{s}_{u},h,\beta)\rfloor\}.$
Proof.
It suffices to combine Proposition 4 and Proposition 11. ∎
We present three theoretical algorithms based on Theorem 5. These algorithms compute $\mathfrak{s}(u)$ and the smallest element of $\operatorname{Argmax}(u)$ which corresponds to $\mathbf{K}_{u}$ (see Section 2). Algorithm 1 describes the most general situation where the computations are done from an upper bound $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$. In Algorithm 2, we consider the case where the upper bound belongs to $\Gamma_{0}^{\rm dec}(u)$. Finally, Algorithm 3 deals with an upper bound $(h,\beta)\in{\Gamma}_{0}^{\rm cst}(u)$. In the algorithms presented, the upper bounds $(h,\beta)$ are supposed to be usefully decreasing for the first case and useful in the two last cases.
1
Input : $u\in\Lambda$ with ${\mathrm{R}}_{u}^{s}\neq\emptyset$ and $(h,\beta)\in\Gamma_{{\rm ev}}^{\rm dec}(u)$ usefully decreasing
Output : $\mathfrak{s}(u)$ and $\mathbf{K}_{u}=\min\operatorname{Argmax}(u)$
2 begin
3       $k=0$; $K=+\infty$, $\operatorname{vmax}=-\infty$, $\operatorname{kmax}=0$
4       while $k<\mathfrak{m}(h,\beta)$ do
5             if $\operatorname{vmax}<u_{k}$ then
6                   $\operatorname{vmax}=u_{k}$
7                   $\operatorname{kmax}=k$
8
9            $k=k+1$
10
11      while $k\leq K$ do
12             if $k\in\mathcal{S}(u,h)$ then
13                   $K=\min\{K,\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor\}$
14                   if $\operatorname{vmax}<u_{k}$ then
15                         $\operatorname{vmax}=u_{k}$
16                         $\operatorname{kmax}=k$
17
18
19            $k=k+1$
20
21      $\mathfrak{s}(u)=\operatorname{vmax}$
22       $\mathbf{K}_{u}=\operatorname{kmax}$
23
24
Algorithm 1 Resolution of Problem (1) with an eventually and usefully decreasing upper bound
In Algorithm 1, to compute a truncation index using $\mathfrak{F}_{u}$, we have to go over $\mathfrak{m}(h,\beta)$ to ensure that we compute an upper bound of $\mathbf{K}^{s}_{u}$. Once $\mathfrak{m}(h,\beta)$ achieved, $\mathfrak{F}_{u}(k,h,\beta)$ is guaranteed to be greater than $\mathbf{K}^{s}_{u}$. However, we do not know for which $k$ greater than $\mathfrak{m}(h,\beta)$, $\mathfrak{F}_{u}(k,h,\beta)$ is minimal. To update the truncation index $K$ permits to minimize the number of iterations. In this general context, $u_{k}$ is supposed to be computable easily. In some specific situations, the computation of $u_{k}$ can be very expensive (e.g. the optimal value of an optimization problem [Adj25]). Finally, the integer maximizer found is $\mathbf{K}_{u}$, the minimum of $\operatorname{Argmax}(u)$, as the update condition at Line 11 of Algorithm 1 is $\operatorname{vmax}<u_{k}$. To get $\mathbf{K}^{s}_{u}$, it suffices to replace $\operatorname{vmax}<u_{k}$ by $\operatorname{vmax}\leq u_{k}$.
1
Input : $u\in\Lambda$ with ${\mathrm{R}}_{u}^{s}\neq\emptyset$ and $(h,\beta)\in\Gamma_{0}^{\rm dec}(u)$ useful
Output : $\mathfrak{s}(u)$ and $\mathbf{K}_{u}=\min\operatorname{Argmax}(u)$
2 begin
3       $k=0$; $K=+\infty$, $\operatorname{vmax}=-\infty$, $\operatorname{kmax}=0$
4       while $k\leq K$ do
5             if $k\in\mathcal{S}(u,h)$ then
6                   $K=\min\{K,\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor\}$
7                   if $\operatorname{vmax}<u_{k}$ then
8                         $\operatorname{vmax}=u_{k}$
9                         $\operatorname{kmax}=k$
10
11
12            $k=k+1$
13
14      $\mathfrak{s}(u)=\operatorname{vmax}$
15       $\mathbf{K}_{u}=\operatorname{kmax}$
16
17
Algorithm 2 Resolution of Problem (1) with a useful decreasing upper bound
In Algorithm 2, we can start computing a truncation index from $\mathfrak{F}_{u}$ from $k=0$ as $\mathfrak{F}_{u}(k,h,\beta)$ is guaranteed to be greater than $\mathbf{K}^{s}_{u}$ since $(h,\beta)$ is decreasing. However, again, we do not know for which index $k$, $\mathfrak{F}_{u}(k,h,\beta)$ is minimal.
1
Input : $u\in\Lambda$ with ${\mathrm{R}}_{u}^{s}\neq\emptyset$ and $(h,\beta)\in{\Gamma}_{0}^{\rm cst}(u)$ useful
Output : $\mathfrak{s}(u)$ and $\mathbf{K}_{u}=\min\operatorname{Argmax}(u)$
2 begin
3       $k=0$; $K=+\infty$, $\operatorname{vmax}=-\infty$, $\operatorname{kmax}=0$
4       while $k\leq K$ do
5             if $k\in\mathcal{S}(u,h)$ then
6                   if $\operatorname{vmax}<u_{k}$ then
7                         $K=\lfloor\mathfrak{F}_{u}(k,h,\beta)\rfloor$
8                         $\operatorname{vmax}=u_{k}$
9                         $\operatorname{kmax}=k$
10
11
12            $k=k+1$
13
14      $\mathfrak{s}(u)=\operatorname{vmax}$
15       $\mathbf{K}_{u}=\operatorname{kmax}$
16
17
Algorithm 3 Resolution of Problem (1) with a useful constant upper bound
Finally, in Algorithm 3, when the upper bound $(h,\beta)\in{\Gamma}_{0}^{\rm cst}(u)$, we can exploit the second statement of Corollary 3 and thus update the truncation index $K$ only when $\operatorname{vmax}$ is updated. The third statement of Proposition 11 proves that the stopping integer cannot be updated once reached $\mathbf{K}_{u}$.
4 Applications
4.1 A simple case
We start with a very simple example to illustrate our approach. Let us consider, for $a\in\mathbb{N}^{*}$, the sequence $x:=(a^{n}(n!)^{-1})_{n\in\mathbb{N}}$. As presented in Problem 1, we are interested in the term $x_{n}$ which maximizes the sequence $x$ and an maximizer index. We apply our techniques to compute them. It is evident that alternative approaches (the study of the variation of the sequence) can solve the problem in a easier way. We can rewrite, for all $n\in\mathbb{N}$, the terms $x_{n}$ as follows:
$x_{n}=\dfrac{a^{n}}{n!}=\left(\dfrac{a}{a+1}\right)^{n}\dfrac{(a+1)^{n}}{n!}.$
Hence, we have for all $n\in\mathbb{N}$:
$x_{n}=h_{n}(\beta^{n})\text{ where }h_{n}:[0,1]\ni t\to t\dfrac{(a+1)^{n}}{n!}\text{ and }\beta=\dfrac{a}{a+1}$
It is straightforward to see that $h=(h_{n})_{n\in\mathbb{N}}$ is eventually decreasing and $\mathfrak{m}(h)=a$. Moreover, as for all $n\in\mathbb{N}$, $h_{n}(0)=0$, we have $\mathcal{S}(x,h)=\mathbb{N}$. Then following Algorithm 1, we can compute $\mathfrak{F}_{x}(k,h,\{\beta\}_{n\in\mathbb{N}})$ from $k=a$. Then we have:
$\mathfrak{F}_{x}(a,h,\{\beta\}_{n\in\mathbb{N}})=\dfrac{\ln\left(h_{a}^{-1}(x_{a})\right)}{\ln(\beta)}=\dfrac{\ln\left(\dfrac{a^{a}}{a!}\times\dfrac{a!}{(a+1)^{a}}\right)}{\ln\left(\dfrac{a}{a+1}\right)}=\dfrac{a\ln\left(\dfrac{a}{a+1}\right)}{\ln\left(\dfrac{a}{a+1}\right)}=a.$
We conclude that we have to compare the $a+1$ first terms of the sequence $x$ to find $\mathfrak{s}(x)$ and a integer maximizer of $x$. Using alternative approaches, we know that the maximal term of the sequence $x$ is achived at $k=a$.
This example shows why we allow sequences of functions in upper bounds of the analyzed sequence whereas Theorem 1 suggests to use a constant sequence determined by an affine upper bound function. For example, defining $g:t\to t(a+1)^{a}$, we have for all $n\in\mathbb{N}$, $x_{n}\leq g(\beta^{n})$. However, as $h_{n}\leq g$ for all $n\in\mathbb{N}$, we have from Corollary 2 $\mathfrak{F}_{u}(a,h,\{\beta\})\leq\mathfrak{F}_{u}(a,\{g\},\{\beta\})$. The main advantage of using a constant sequence is that we can compute a truncation index from $n=0$ and $\mathfrak{F}_{u}(n,\{g\},\{\beta\})$ is an upper bound of $\mathbf{K}^{s}_{x}$ for all $n\in\mathbb{N}$. More precisely, we have for all $n\in\mathbb{N}$:
$\mathfrak{F}_{x}(n,\{g\},\{\beta\})=\dfrac{\ln(g^{-1}(x_{n})}{\ln(\beta)}=\dfrac{\ln\left(\dfrac{a^{n}}{a!(a+1)^{a}}\right)}{\ln\left(\dfrac{a}{a+1}\right)}=\dfrac{\ln\left(\dfrac{a^{n-a}}{a!}\right)}{\ln\left(\dfrac{a}{a+1}\right)}+a$
We observe that the fifth statement of Theorem 3 holds that is as $x$ is increasing on $[0,a]\cap\mathbb{N}$, $\mathfrak{F}_{x}(n,\{g\},\{\beta\})$ is decreasing on $[0,a]\cap\mathbb{N}$. The smallest value of $\mathfrak{F}_{x}(n,\{g\},\{\beta\})$ is achieved at $n=a$ and
$\mathfrak{F}_{x}(a,\{g\},\{\beta\})=-\dfrac{\ln(a!)}{\ln\left(\dfrac{a}{a+1}\right)}+a.$
This leads to, for example, $\lfloor\mathfrak{F}_{x}(2,\{g\},\{\beta\})\rfloor=3$, $\lfloor\mathfrak{F}_{x}(5,\{g\},\{\beta\})\rfloor=31$ and $\lfloor\mathfrak{F}_{x}(10,\{g\},\{\beta\})\rfloor=168$. The overapproximation gap totally explodes. This justifies to select the tightest possible upper bound of the analyzed sequence even if the upper bound is not a constant sequence.
4.2 Logistic Sequences
Let consider logistic sequences parameterized by $r\in(0,4]$ starting at $y_{0}\in(0,1)$. More precisely, we have for all $n\in\mathbb{N}$:
$y_{n+1}:=ry_{n}(1-y_{n})$
We can exploit closed forms for $r\in(0,1)$. For $r\in(0,1)$, from [CB19, Lemma A.6], we have at hand the upper-bound for all $n\in\mathbb{N}$, $y_{n}\leq y_{0}(r^{-n}+y_{0}n)^{-1}$. We then have for all $n\in\mathbb{N}$, $y_{n}\leq h_{n}(\beta^{n})$ where $h_{n}:(0,1]\ni t\mapsto y_{0}(t^{-1}+ny_{0})^{-1}$ and $h_{n}(0)=0$ and $\beta=r$. Then, for all $n\in\mathbb{N}$, $h_{n}$ belongs to $\Omega([0,1])$ and the sequence $(h_{n})_{n}$ is decreasing. Moreover, $y_{0}=h_{0}(1)$ and we can then apply the third statement of Proposition 6 to conclude that $\mathfrak{s}(y)=y_{0}$ and $\operatorname{Argmax}(y)=\{0\}$.
In the general setting $r\in(0,4]$, we can always find an upper-bound $(h(\beta^{n}))_{n}$ for $y$. Indeed, as $t\mapsto rt(1-t)$ is bounded by $r/4$ on $[0,1]$. Then we have $y_{n}\leq r/4+\beta^{n}$ for all $\beta\in[0,1)$. Hence, we have, $y_{n}\leq h(\beta^{n})$ where $h:x\mapsto r/4+x$. However, this function $h$ is not useful since $y_{n}\leq h(0)$ for all $n\in\mathbb{N}$. The key difficulty is to find a useful functional upper bound of the sequence.
4.3 The sequence of the ratio of two consecutive terms of a Fibonacci sequence
In this example, we consider the ratio of some Fibonacci sequences starting from $u_{0}\geq 0$ and $u_{1}>u_{0}\phi$ (for simplicity) where $\phi=(1+\sqrt{5})/2$ (the golden number). By definition of Fibonacci sequences, we have for all $n\in\mathbb{N}$, $u_{n+2}=u_{n+1}+u_{n}$. It is well-known that for all $n\in\mathbb{N}$, $u_{n}=A\phi^{n}+B(-\phi)^{-n}$ where $A=\frac{u_{0}+u_{1}\phi}{1+\phi^{2}}$ and $B=\frac{(u_{0}\phi-u_{1})\phi}{1+\phi^{2}}$. From, the assumption on $u_{0}$ and $u_{1}$, we have $A>0$ and $B<0$ but $u_{n}>0$ for all $n\in\mathbb{N}^{*}$. Let us define:
$\omega_{0}:=\left\{\begin{array}[]{lr}u_{1}/u_{0}&\text{ if }u_{0}>0\\ 0&\text{ if }u_{0}=0\end{array}\right.\text{ and for all }n\in\mathbb{N}^{*},\omega_{n}:=\dfrac{u_{n+1}}{u_{n}}\enspace.$
We have for all $n\in\mathbb{N}^{*}$:
$\omega_{n}=\phi\left(1+\frac{B\phi^{-2}(-1)^{n+1}-B}{A\phi^{2n}+B(-1)^{n}}\right)\enspace.$
This proves that $\boldsymbol{\ell}(\omega)=\lim_{n\to+\infty}\omega_{n}=\phi$. This also proves that for all $n$ odd, $\omega_{n}<\phi$ and for all $n$ even and not null, $\omega_{n}>\phi$. Now as $B<0$, we have for all $n\in\mathbb{N}^{*}$:
$\omega_{n}\leq h(\beta^{n})\text{ where }h:x\to\left\{\begin{array}[]{lr}\phi\left(1-\dfrac{B\phi^{-2}+B}{Ax^{-1}+B}\right)&\text{ if }x\neq 0\\ \\ \phi&\text{ if }x=0\end{array}\right.\text{ and }\beta=\phi^{-2}$
with the equality when $n$ is even.
The function $h$ is strictly increasing and continuous on $[0,1]$ and $\beta\in(0,1)$ thus ($(\{h\}_{n},\{\beta\}_{n})\in{\Gamma}_{0}^{\rm cst}(\omega)$. When $n\in\mathbb{N}^{*}$ is even and thus the equality holds, we have $h^{-1}(\omega_{n})=(\phi^{-2})^{n}$. For the case where $n=0$, we also have $\omega_{0}\leq h(1)$. If $u_{0}$ is null, 00 does not belong to $\mathcal{S}(\omega,h)$. If $u_{0}$ is strictly positive, the equality $u_{0}=h(1)$ holds and we can conclude from the third statement of Proposition 6 that $\mathfrak{s}(\omega)=\omega_{0}$ and $\operatorname{Argmax}(\omega)=\{0\}$.
Now we suppose that $u_{0}=0$. We can evalute $\mathfrak{F}_{\omega}(k,\{h\}_{n\in\mathbb{N}},\{\beta\}_{n\in\mathbb{N}})$ for all $k$ such that $\omega_{k}>h(0)=\phi$. Actually, $\mathcal{S}(\omega,h)=\{2k:k\in\mathbb{N}^{*}\}$. For all $k\in\mathbb{N}^{*}$,
$\mathfrak{F}_{\omega}(2k,\{h\}_{n\in\mathbb{N}},\{\beta\}_{n\in\mathbb{N}})=\dfrac{\ln(h^{-1}(\omega_{2k}))}{\ln(\phi^{-2})}=(2k)\dfrac{\ln(\phi^{-2})}{\ln(\phi^{-2})}.$
We conclude that $\mathbf{K}^{s}_{\omega}\leq 2$ and since $\omega_{0}$ and $\omega_{1}$ are smaller than $\phi$, $\mathbf{K}^{s}_{\omega}=2$.
4.4 Syracuse sequences
Let consider a (accelerated) Syracuse iteration starting at $N\in\mathbb{N}^{*}$. Formally, we define the following sequence:
$y_{0}:=N\text{ and for all }n\in\mathbb{N},\ y_{n+1}:=\left\{\begin{array}[]{lr}\dfrac{y_{n}}{2}&\text{ if }y_{n}\text{ is even}\\ \\ \dfrac{3y_{n}+1}{2}&\text{ if }y_{n}\text{ is odd}\end{array}\right.$ (10)
The Collatz conjecture asserts that for all $N\in\mathbb{N}^{*}$, the sequence $(y_{n})_{n}$ defined according to (10) admits an index $k\in\mathbb{N}$ such that $y_{k}=1$. The Collatz conjecture trivially holds when we start with an integer $N\in\{1,2,3,4\}$.
Theorem 6.
The Collatz conjecture holds for all integers $N\geq 5$ if and only if there exists $(a,b,c)\in\mathbb{R}_{+}\times(0,1)\times(4,5]$ such that $y_{n}\leq ab^{n}+c$ for all $n\in\mathbb{N}$.
Proof.
If the Collatz conjecture holds then $\boldsymbol{\ell}(y)=4$ as $(1,4,2,1)$ is a cycle of the sequence. Since we suppose that $N\geq 5$, ${\mathrm{R}}_{u}^{s}$ is nonempty and we can follow the proof of Lemma 1 and choose $c\in(4,5]\subset(\boldsymbol{\ell}(y),\mathfrak{s}(y))$. Again following the proof of Th 1, we deduce suitable $a$ and $b$ to obtain the inequality.
Suppose that the inequality holds with $(a,b,c)\in\mathbb{R}_{+}\times(0,1)\times(4,5]$. As $c\in(4,5]$, then $\boldsymbol{\ell}(y)\leq c\leq 5$. As $y$ is a sequence of strictly positive integers, it implies that there exists $M\in\mathbb{N}$ such that $y_{n}\in\{1,2,3,4,5\}$ for all $n\geq M$. The values 3 and 5 are not possible. Indeed, if $y_{n}=3$, $y_{n+1}=5$ and $y_{n+2}=8$ which is not less than 5. ∎
4.5 Linear Discrete-time dynamical systems
4.5.1 General case
Let consider a $d\times d$ matrix $A$ and the trajectories $x_{k+1}=Ax_{k}$ starting from $x_{0}\in\mathbb{R}^{d}$. We suppose that $A$ is stable i.e. its spectral radius is strictly less than 1.
Recall that all positive definite $d\times d$ matrices $Q$ define a norm on $\mathbb{R}^{d}$ as follows $\left\|x\right\|_{Q}=\sqrt{x^{\intercal}Qx}$. From this, we can define an operator norm for $A$ associated to $Q$ as follows $\left\|A\right\|_{Q}^{2}=\max_{x\neq 0}\left\|Ax\right\|_{Q}/\left\|x\right\|_{Q}$. This norm is submultiplicative i.e. $\left\|BC\right\|_{Q}\leq\left\|B\right\|_{Q}\left\|C\right\|_{Q}$ and thus implies that $\left\|A^{k}\right\|_{Q}\leq\left\|A\right\|_{Q}^{k}$. The fact that $A$ is stable is equivalent to that for all matrix norms $\left\|A^{k}\right\|$ tends to 0 as $k$ goes to +$\infty$+ ∞.
In [Shc17], the authors are interested in the maximum of the Euclidean matrix norm of the powers of $A$ i.e. $\max_{k\in\mathbb{N}}\|A^{k}\|_{2}$ recalling that $\left\|A\right\|_{2}^{2}=\max_{x\neq 0}\left\|Ax\right\|_{2}/\left\|x\right\|_{2}$. Following [Adj25], we can use solutions of the discrete Lyapunov equation i.e. a matrix from $\mathcal{L}_{A}:=\{P\succ 0:P-A^{\intercal}PA\succ 0\}$ where the symbol $R\succ 0$ means that $R$ is positive definite. Now taking $P\in\mathcal{L}_{A}$, we can define the operator norm $\left\|A\right\|_{P}$ as $P$ is positive definite. The main advantage of using solutions of the discrete Lyapunov equation is to ensure that $\left\|A\right\|_{P}<1$. Moreover, let $k\in\mathbb{N}$ and let $x\in\mathbb{R}^{d}$ such that $A^{k}x\neq 0$. Then:
$\begin{array}[]{lcl}\displaystyle{\dfrac{x^{\intercal}{A^{k}}^{\intercal}A^{k}x}{x^{\intercal}x}}&=&\displaystyle{\dfrac{x^{\intercal}{A^{k}}^{\intercal}A^{k}x}{x^{\intercal}{A^{k}}^{\intercal}P{A^{k}}x}\dfrac{x^{\intercal}{A^{k}}^{\intercal}PA^{k}x}{x^{\intercal}Px}\dfrac{x^{\intercal}Px}{x^{\intercal}x}}\\ &\leq&\displaystyle{\max_{y\neq 0}\dfrac{y^{\intercal}y}{y^{\intercal}Py}\left\|A^{k}\right\|_{P}^{2}\lambda_{\rm max}(P)}\\ &=&\displaystyle{\dfrac{1}{\displaystyle{\min_{y\neq 0}\frac{y^{\intercal}Py}{y^{\intercal}y}}}\left\|A^{k}\right\|_{P}\lambda_{\rm max}(P)}\\ &=&\displaystyle{\dfrac{\lambda_{\rm max}(P)}{\lambda_{\rm min}(P)}\left\|A^{k}\right\|_{P}^{2}}\leq\displaystyle{\dfrac{\lambda_{\rm max}(P)}{\lambda_{\rm min}(P)}\left\|A\right\|_{P}^{2k}}\end{array}$
If $A^{k}$ is not invertible, the inequality also holds for all non-zero $x\in\mathbb{R}^{d}$ in the kernel of $A^{k}$. Finally taking the maximum over all non-zero $x\in\mathbb{R}^{d}$, we get for all $k\in\mathbb{N}$
$\left\|A^{k}\right\|_{2}^{2}\leq h_{P}(\beta_{P}^{k})\text{ where }h_{P}:t\to\dfrac{\lambda_{\rm max}(P)}{\lambda_{\rm min}(P)}t\text{ and }\beta_{P}=\left\|A\right\|_{P}^{2}\enspace.$
Finally, let for all $k\in\mathbb{N}$, $z_{k}:=\left\|A^{k}\right\|_{2}^{2}$ and define $z=(z_{k})_{k\in\mathbb{N}}$ we have $(h_{P},\beta_{P})\in{\Gamma}_{0}^{\rm cst}(z)$. We can apply our method to compute the maximal value of the sequence $z$. As the terms $z_{k}$ are strictly positive and for all $P\in\mathcal{L}_{A}$, $h_{P}(0)=0$, we have $\mathcal{S}(z,h_{P})=\mathbb{N}$. Moreover, we have:
$\mathfrak{F}_{z}(k,\{h_{P}\}_{n\in\mathbb{N}},\{\beta_{P}\}_{n\in\mathbb{N}})=\dfrac{\ln\left(h_{P}^{-1}(z_{k})\right)}{\ln(\beta_{P})}=\dfrac{\ln\left(\dfrac{\lambda_{\rm min}(P)}{\lambda_{\rm max}(P)}\left\|A^{k}\right\|_{2}^{2}\right)}{2\ln(\left\|A\right\|_{P}}.$
which provides upper-bounds of the greatest maximizer of $z$.
4.5.2 Numerical example
Let us consider one example developed in [APS18]. Let us define $U$ as the $d\times d$ matrix such that $U_{1d}=1$ and $U_{ij}=0$ otherwise and $\mathrm{Id}$ the identity matrix (of size $d\times d$). Now let $\lambda$ be in $(-1,1)$ and let us define we define the matrix:
$A_{\lambda}:=\begin{pmatrix}\lambda&0&\ldots&1\\ 0&\ddots&\ddots&\vdots\\ \vdots&\vdots&\lambda&0\\ 0&\ldots&0&\lambda\end{pmatrix}=\lambda\mathrm{Id}+U$
Then, as $U^{k}=0$ for all $k\geq 2$, we have for all $k\in\mathbb{N}$ $A_{\lambda}^{k}=(\lambda Id+U)^{k}=\lambda^{k}\mathrm{Id}+k\lambda^{k-1}U$. Moreover, $\left\|A_{\lambda}^{k}\right\|_{2}^{2}=\lambda_{\rm max}((A_{\lambda}^{k})^{\intercal}A_{\lambda}^{k})=\lambda^{2k-2}(\lambda^{2}+0.5k^{2}+0.5k\sqrt{4\lambda^{2}+k^{2}})$. The goal is thus to compute $\max_{k\in\mathbb{N}}\left\|A_{\lambda}^{k}\right\|_{2}^{2}$ by giving an upper-bound of the (greatest) maximizer of $\max_{k\in\mathbb{N}}\left\|A_{\lambda}^{k}\right\|_{2}^{2}$.
For the class of matrices, $\mathcal{L}_{A_{\lambda}}$ contains particular diagonal matrices. Indeed, $P\in\mathcal{L}_{A_{\lambda}}$ if and only if $P\succ 0$ and :
$P-(A_{\lambda})^{\intercal}P(A_{\lambda})=P(1-\lambda^{2})-\lambda(U^{\intercal}P+PU)+U^{\intercal}PU\succ 0$
To simplify, we consider diagonal matrices $P$ which leads to consider positive definite diagonal matrices $P$ such that that the matrix;
$P-A_{\lambda}^{\intercal}PA_{\lambda}=\begin{pmatrix}(1-\lambda^{2})p_{1,1}&0&\ldots&-\lambda p_{11}\\ 0&(1-\lambda^{2})p_{2,2}&\ddots&\vdots\\ \vdots&\vdots&\ddots&0\\ -\lambda p_{11}&0&\ldots&(1-\lambda^{2})p_{d,d}-p_{1,1}\end{pmatrix}$
is positive definite.
Then, we can take $p_{1,1}=p_{2,2}=\cdots=p_{d-1,d-1}=1$ and in this case $P-A_{\lambda}^{\intercal}PA_{\lambda}\succ 0$ if and only if its determinant is strictly positive. As $\det(P-A_{\lambda}^{\intercal}PA_{\lambda})=(1-\lambda^{2})^{d-2}(p_{d,d}(1-\lambda^{2})^{2}-1)$, we conclude that $p_{d,d}$ must be strictly greater than $(1-\lambda^{2})^{-2}$. Finally, we define $P_{q}:=\operatorname{Diag}(1,\cdots,1,q)$ where $\operatorname{Diag}(v)$ is the diagonal matrix for which the diagonal is the vector $v$ and $q\in\left((1-\lambda^{2})^{-2},+\infty\right)$.
For those matrices $P_{q}$, we have:
$\left\|A_{\lambda}\right\|_{P_{q}}^{2}=\dfrac{1+2q\lambda^{2}+\sqrt{1+4\lambda^{2}q}}{2q}$
Finally,
$\mathfrak{F}_{z}(k,\{h_{P_{q}}\}_{n\in\mathbb{N}},\{\beta_{P_{q}}\}_{n\in\mathbb{N}})=\dfrac{\ln\left(\dfrac{\lambda^{2k-2}(\lambda^{2}+0.5k^{2}+0.5k\sqrt{4\lambda^{2}+k^{2}}}{q}\right)}{\ln\left(\dfrac{1+2q\lambda^{2}+\sqrt{1+4\lambda^{2}q}}{2q}\right)}.$
For our experiments, we have chosen $q=2(1-\lambda^{2})^{-2}$. Table 1 shows the computations for several values of $\lambda$.
$\mathbf{K}^{s}_{z}$ ${\max_{k\in\mathbb{N}}z_{k}}$ $\mathfrak{F}_{z}(\mathbf{K}^{s}_{z})$
$\lambda=0.1$ 0 1 1
$\lambda=0.25$ 0 1 1
$\lambda=0.5$ 1 1.4572 2
$\lambda=0.75$ 3 3.1937 7
$\lambda=0.9$ 9 15.3082 20
$\lambda=0.99$ 99 1367.27 221
$\lambda=0.999$ 999 135471 2229
$\lambda=0.99995$ 19999 54136820.5 44617
Table 1: Numerical results
4.5.3 Going beyond
We could consider as suggested in this paper non constant sequences $(P_{k})_{k\in\mathbb{N}}$ in $\mathcal{L}_{A_{\lambda}}$. Following Corollary 1, we must impose the decreasing condition that is in this case, for all $k\in\mathbb{N}$:
$\dfrac{\lambda_{\rm max}(P_{k+1})}{\lambda_{\rm min}(P_{k+1})}\leq\dfrac{\lambda_{\rm max}(P_{k})}{\lambda_{\rm min}(P_{k})}\text{ and }\left\|A_{\lambda}\right\|_{P_{k+1}}\leq\left\|A_{\lambda}\right\|_{P_{k}}$
Note that $\left\|A_{\lambda}\right\|_{P}$ is lower bounded by the spectral radius of $A_{\lambda}$ which is equal to $|\lambda|$. This is also suggested in [Adj25] by considering a minimization problem over $\mathcal{L}_{A_{\lambda}}$:
$\begin{array}[]{rcl}\operatorname{Min}&&F_{k}(P)\\ &\text{s.t.}&P\in\mathcal{L}_{A_{\lambda}}\end{array}\text{ where }F_{k}(P):=\dfrac{\ln\left(\dfrac{\lambda_{\rm min}(P)}{\lambda_{\rm max}(P)}\left\|A^{k}\right\|_{2}^{2}\right)}{2\ln(\left\|A_{\lambda}\right\|_{P}}$
In this case, monotonicity is a stronger condition (from the integer $N$, if any, such that the sequence $(\left\|A_{\lambda}^{k}\right\|_{2})_{k}$ is decreasing ) compared with the condition $P_{k}$ minimizes $F_{k}$ over $\mathcal{L}_{A_{\lambda}}$.